### v1

# A Hypernetwork-Based Approach to KAN Representation of Audio Signals

Abstract
Implicit neural representations (INR) have gained prominence for efficiently encoding multimedia data, yet their applications in audio signals remain limited. This study introduces the Kolmogorov-Arnold Network (KAN), a novel architecture using learnable activation functions, as an effective INR model for audio representation. KAN demonstrates superior perceptual performance over previous INRs, achieving the lowest Log-Spectral Distance of 1.29 and the highest Perceptual Evaluation of Speech Quality of 3.57 for 1.5 s audio. To extend KAN’s utility, we propose FewSound [^1] The source code can be accessed at https://github.com/gmum/fewsound.git, a hypernetwork-based architecture that enhances INR parameter updates. FewSound outperforms the state-of-the-art HyperSound, with a 33.3% improvement in MSE and 60.87% in SI-SNR. These results show KAN as a robust and adaptable audio representation with the potential for scalability and integration into various hypernetwork frameworks.
representation learning, audio representation, INR, KAN
1 Introduction
Implicit neural representations (INRs) represent discrete data in neural network weights. INR finds many applications in the case of image (Klocek et al., 2019), video (Chen et al., 2022), or 3D objects (Spurek et al., 2020). In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data. HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes on raw input sound and produces weight for INR-based sound representation. INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes. Siamese SIREN (Lanzendörfer & Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
The recently proposed KAN (Kolmogorov-Arnold Network) (Liu et al., 2024a) replaces traditional fully connected layers with activation functions. While KAN shows potential, it faces limitations with high-dimensional data. However, its effectiveness in low-dimensional settings and successful application to complex function regression tasks make it a promising candidate for audio INR.
This paper demonstrates that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm. For single audio INR, KAN trains more rapidly and achieves better results. Setting up the conditions for a KAN is not immediately straightforward. We show that hypernetworks can generate KANs. In practice, we can directly produce KAN parameters or use updates for the universal KAN-based architecture. Therefore, we propose two methods. The first is a classical HyperSound (Szatkowski et al., 2023) with a KAN target network. The second uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR, which is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017). Since in a few-shot approach, there does not exist a model dedicated to audio signals, we introduce FewSound. FewSound takes raw sound on input and uses hypernetwork to generate updates to universal weights of INR dedicated to encoding sound. In our work, we examine six approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a). In the model, we simultaneously train the hypernetwork weights and the universal weights. Thus, two settings show that KAN can be effectively modeled by hypernetworks.
Our study demonstrates that KAN model can be viewed as INR of sound. Such models yield favorable outcomes when dealing with individual audio signals and excel in large Hypernetwork-based models trained on large datasets.
Key contributions of our research include:
  • showing KAN’s efficacy for continuous sound representation,
  • comparing KAN represenation with other INRs,
  • showing that KAN can be produced via hypernetworks,
  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN network and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
  • analysis of the parameters’s influence on the efficacy of the proposed method.
2 Related Work
To our knowledge, there are no specialized INR architectures focused on sound. Research papers on INR (Sitzmann et al., 2020; Tancik et al., 2020; Saragadam et al., 2023; Liu et al., 2024b; Müller et al., 2022) typically evaluate models using images, videos, and 3D objects as benchmarks. Although experiments involving sounds do exist, they remain relatively unexplored. In this section, we show that KAN effectively models single-audio in an efficient manner.
INR
Implicit neural representations (INR) are data approximations based on neural networks relying on coordinate inputs. They enable discrete information such as pixels (Klocek et al., 2019; Chen et al., 2022), cloud points (Spurek et al., 2020), or audio magnitude information (Szatkowski et al., 2023) to be represented as continuous functions. The separation of the representation from spatial resolution makes the representation more memory efficient. Such information is limited by the networks’ architecture instead of the grid resolution. INRs are used in many signal processing problems, including 3D reconstruction (Mildenhall et al., 2020), super-resolution (Chen et al., 2022), compression (Strümpler et al., 2022; Lanzendörfer & Wattenhofer, 2023) or audio scene representation (Su et al., 2022).
Initially, the INRs used ReLU-based MLPs. However, due to the ReLU’s second derivative being zero, they cannot capture the fine details of the information located in higher-order derivatives leading to worse performance. To counteract this, methods such as SIREN (Sitzmann et al., 2020) have been developed, which use a periodic activation function the sine in place of ReLU. The i-th layer of the network is defined as: $\phi_{i}(x_{i})=\sin(\omega_{i}W_{i}^{T}x_{i}+b_{i}),$ where $x_{i}$ is layer’s input, while $W_{i}$ and $b_{i}$ are, respectively, weights and biases. The use of the sine function enables SIREN to capture the details of the processed signals to a much greater degree than the widely used activation functions. Another example of a network architecture using a specialized activation function is WIRE (wavelet implicit neural representation) (Saragadam et al., 2023). Inspired by harmonic analysis, WIRE uses complex Gabor wavelet $\psi$ as its nonlinearity activation: $\psi(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$.
INRs can suffer from a phenomenon called spectral bias (Rahaman et al., 2019) – the preference of MLPs to learn low-frequency functions and ignore high-frequency noise, which can lead to performance degradation when high-frequency components of the input carry meaningful information (Gorji et al., 2023). A common way to overcome spectral bias is by introducing high-frequency embeddings (Mildenhall et al., 2020; Tancik et al., 2020) that modify the space over which the MLP operates, altering the frequency of its output function. FINER (Liu et al., 2024b) is an MLP-based architecture designed to flexibly tune the spectral bias. It uses a variable-periodic activation $a(x)=sin(\omega_{0}\lvert x+1\rvert x)$ and a specialized bias initialization schema in its first layer.
NERF (Neural Radiance Field) (Mildenhall et al., 2020) is a type of network originally designed for scene representation. Based on the 5D input of spatial location $(x,\,y,\,z)$ and viewing direction ($\theta,\,\phi$), the MLP network outputs color c = $(r,\,g,\,b)$ and volume density ($\sigma$). This allows the generation of 3D reconstructions of scenes from a set of sparse views.
Sound representation
The audio encoders we rely on are classified into two main types: neural speech codecs and those employed in automatic speech recognition (ASR) or ASR-adjacent solutions.
SoundStream (Zeghidour et al., 2021) is a popular neural audio codec based on the encoder-decoder architecture. In our experiments we used only the encoder part composed of 1D convolution followed by four encoder blocks with residual units and the final 1D convolution.
SNAC (Siuzdak, 2024) uses the encoder based on the hierarchical tokens like the previous approach. The distinction is that the coarse tokens are sampled less frequently, resulting in covering broader time windows. SNAC not only decreases the bitrate required for reconstruction, but also provides useful information in language modeling tasks.
Whisper (Radford et al., 2023) is a Transformer-based (Vaswani et al., 2017) ASR system, trained on 680,000 hours of data, achieving state-of-the-art speech recognition results. It was successfully adapted to various other tasks including audio-language models (Chu et al., 2023), voice conversion (Guo et al., 2023), text-to-speech (Collabora, 2024) and audio DeepFake detection (Kawa et al., 2023). In our experiments, we use only Whisper’s encoder.
Spectrogram Encoder is based on VoiceBlock (O’Reilly et al., 2022) — originally developed as a speech-to-speech solution for adversarial examples against ASR models. The encoder is based on the linear-spectrogram input that extracts a feature vector composed of fundamental frequency, aperiodicity and loudness of the signal.
KAN (Kolmogorov-Arnold Network)
KANs (Liu et al., 2024a) is a type of neural network architecture used as an alternative to MLPs, instead of using fixed activation functions. KANs utilize learnable activation functions placed on edges rather than nodes. KAN’s learnable functions are parametrized as a spline and serve as ”weights”, whereas the nodes only sum the incoming signals. This is described by the following formula:
$f(x)=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,\,p}(x_{p})\right)~{},$
where $\Phi_{q}$ is a KAN layer, $\phi_{q,\,p}$ is a single activation function and $x_{p}$ is input feature. The authors also introduced an architecture of LANs (learnable activation networks) – the only difference between them and MLP lies in the learning aspect of the activation functions. The authors show that a network can be used in implicit neural representation tasks.
Audio INR
In the realm of acoustic data, INR transforms a one-dimensional temporal space into the corresponding amplitude values. Therefore, our INR is depicted as:
$f_{\theta}:\mathbb{R}\to\mathbb{R}~{},$
where $f_{\theta}(t)$ denotes the amplitude of the sound over time.
We consider six variants of INRs: NeRF, SIREN, RFF, FINER, WIRE and KAN. We use NeRF to refer to a network with positional encoding, followed by a MLP with ReLU activation. The positional encoding, described by $\gamma(t)=\left[\sin(2^{0}\pi t),\cos(2^{0}\pi t),\ldots,\sin(2^{L-1}\pi t),\cos(2^{L-1}\pi t)\right]$ is taken from the original NeRF (Mildenhall et al., 2020).
SIREN (Sitzmann et al., 2020) is a MLP with sine activation: $a(x)=\sin(\omega_{0}x)$, where $\omega_{0}$ is a hyperparameter.
RFF (Random Fourier Features) (Tancik et al., 2020) is used to refer to a multilayer perceptron with ReLU activation and input embedding defined as $\gamma(t)=\left[\cos(2\pi\mathbf{B}t),\sin(2\pi\mathbf{B}t)\right],$ where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is a random projection matrix, sampled from $\mathcal{N}(0,\sigma^{2})$. In our case $d$, the dimensionality of the input, is always equal to 1, while $\sigma$ and $m$ are hyperparameters.
WIRE (Saragadam et al., 2023) is a MLP with complex Gabor wavelet activation $a(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$, where $\omega_{0}$ and $s_{0}$ are hyperparameters.
FINER (Liu et al., 2024b) is a MLP with activation function defined as $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$, where $\omega_{0}$ is a hyperparameter. The bias of its first layer is initialized by sampling from $\mathcal{U}(-k,k)$, where $k$ is a hyperparameter.
Our Kolmogorov-Arnold network implementation consists of the same positional encoding as in NeRF, followed by KAN layers. Each KAN layer can be described by $\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}spline(x),$ where $w_{b}$ and $w_{s}$ are matrices of learnable weights, and $spline(x)$ is a linear combination of B-splines such that $spline(x)=\sum_{i}{c_{i}B_{i}(x)},$ with learnable $c_{i}$s. It is parametrized by spline order – the degree of the used spline, and grid size – the number of knots defining the spline.
3 KAN in hypernetwork-based architectures
In this section, we show how KAN works with hypernetworks. We use two main approaches. First, we use the autoencoder-based solution HyperSound (Szatkowski et al., 2023), where we add KAN as the target model. We also examine KAN in a few-shot setting introduced by (Wang et al., 2020). To examine such a model, we propose FewSound, which is a realization of (Wang et al., 2020) in the case of KAN and hypernetworks.
Hypernetwork
In the foundational study (Ha et al., 2016) hyper-networks are described as neural models that produce weights for another network tasked with solving a specific problem. The researchers aim to decrease the number of trainable parameters by creating a hyper-network with fewer parameters than the target network. Drawing a parallel between hypernetworks and generative models, the authors of (Sheikh et al., 2017) employ this technique to generate a variety of target networks that approximate the same function.
Figure 1: The overview of FewSound architecture. The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights. The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$. The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
Table 1: Reconstruction metrics for 1.5 s and 3 s audio. We report mean and standard deviation across 110 samples, each of a different speaker of VCTK dataset (Veaux et al., 2013).
1.5 s audio
INR (# params) MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
NERF (33,261) 8.1e-4 $\pm$ 1.3e-3 39.68 $\pm$ 5.67 1.36 $\pm$ 0.18 17.83 $\pm$ 6.10 3.21 $\pm$ 0.48 0.98 $\pm$ 0.04 0.25 $\pm$ 0.11
SIREN (31,191) 4.4e-4 $\pm$ 4.9e-4 40.82 $\pm$ 3.68 1.40 $\pm$ 0.26 19.71 $\pm$ 2.67 2.76 $\pm$ 0.43 0.95 $\pm$ 0.06 0.23 $\pm$ 0.14
FINER (31,191) 1.9e-4 $\pm$ 2.3e-4 44.39 $\pm$ 3.61 2.06 $\pm$ 0.27 23.29 $\pm$ 4.29 2.26 $\pm$ 0.29 0.93 $\pm$ 0.08 0.51 $\pm$ 0.11
WIRE (31,191) 4.8e-4 $\pm$ 6.5e-4 41.91 $\pm$ 5.53 1.63 $\pm$ 0.25 20.77 $\pm$ 5.91 2.79 $\pm$ 0.44 0.96 $\pm$ 0.05 0.16 $\pm$ 0.08
RFF (39,201) 8.7e-4 $\pm$ 8.8e-4 38.63 $\pm$ 4.93 1.56 $\pm$ 0.16 17.81 $\pm$ 5.58 2.94 $\pm$ 0.39 0.98 $\pm$ 0.01 0.19 $\pm$ 0.09
KAN (33,768) 6.7e-4 $\pm$ 2e-3 42.44 $\pm$ 6.72 1.29 $\pm$ 0.22 20.50 $\pm$ 6.94 3.57 $\pm$ 0.42 0.99 $\pm$ 0.02 0.13 $\pm$ 0.08
3 s audio
INR (# params) MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
NERF (33,261) 4.6e-4 $\pm$ 5.0e-4 40.83 $\pm$ 4.50 1.47 $\pm$ 0.16 17.54 $\pm$ 4.29 2.54 $\pm$ 0.45 0.94 $\pm$ 0.05 0.30 $\pm$ 0.14
SIREN (31,191) 5.3e-4 $\pm$ 3.6e-4 39.20 $\pm$ 3.35 1.42 $\pm$ 0.18 14.41 $\pm$ 4.44 2.25 $\pm$ 0.28 0.96 $\pm$ 0.03 0.41 $\pm$ 0.14
FINER (31,191) 1.9e-4 $\pm$ 1.8e-4 43.77 $\pm$ 3.19 2.21 $\pm$ 0.16 19.05 $\pm$ 3.51 1.89 $\pm$ 0.25 0.95 $\pm$ 0.07 0.45 $\pm$ 0.12
WIRE (31,191) 4.3e-4 $\pm$ 3.9e-4 40.30 $\pm$ 3.42 1.52 $\pm$ 0.10 15.45 $\pm$ 4.36 2.33 $\pm$ 0.35 0.95 $\pm$ 0.07 0.33 $\pm$ 0.13
RFF (39,201) 5.3e-4 $\pm$ 3.6e-4 39.20 $\pm$ 3.35 1.42 $\pm$ 0.18 14.41 $\pm$ 4.44 2.25 $\pm$ 0.28 0.96 $\pm$ 0.03 0.40 $\pm$ 0.14
KAN (33,768) 3.8e-4 $\pm$ 4.4e-4 42.28 $\pm$ 5.31 1.34 $\pm$ 0.19 18.85 $\pm$ 5.08 2.89 $\pm$ 0.51 0.97 $\pm$ 0.03 0.33 $\pm$ 0.15
Few-shot learning for INR
A significant limitation of deep learning models is their reliance on extensive datasets. In contrast, humans can learn effectively from just a few examples. A few-shot learning (Wang et al., 2020) approach was proposed to bridge the gap between deep neural network training and human learning capabilities. This strategy treats the data set as consisting of numerous small tasks. Few-shot architectures can be quickly adapted to solve new problems in a single step. The most popular approach to few-shot learning is MAML (Finn et al., 2017). MAML uses a set of universal weights, fine-tuned for specific small tasks through a limited number of gradient updates. Due to the model-agnostic nature of the approach, it can be adapted to various models and corresponding learning algorithms.
The terminology of few-shot learning is often inconsistent due to varying definitions in existing research. For a standardized framework (Chen et al., 2019; Wang et al., 2020). However, such a nomenclature is dedicated to a classification, regression, or generative model task. In the case of the INR-based model, we refer to (Tancik et al., 2021). We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
FewSound
Now we are ready to present FewSound – a model that uses Hypernetworks to update the INR weights dedicated to sound. The core concept of this update mechanism is to use information extracted from raw sound data and universal weights to determine the optimal updates for a particular task. This method adjusts the INR’s parameters for entirely different sounds.
The schema of FewSound is provided in Fig. 1. We aim to reconstruct the training example using the target network. In practice, we run our INR on all steps to reconstruct the input.
Following our previous analysis of INR in the sound reconstruction task, we consider the parameterized function $f_{\theta}$ to be one of the discussed INRs. In addition, we distinguish the trainable encoding network $E(\cdot)$ in our architecture, which transforms raw audio data into low-dimensional representation. We investigated various encoder architectures and report their performance in Tab. 5.
The encoded sound is concatenated with universal weights propagated by the universal weight encoder $\mathbf{G}(\cdot)$ and then to the fully connected hypernetwork $H(\cdot)$. The hypernetwork then modifies the universal weights: $\theta^{\prime}=\theta+\Delta_{\theta}.$ In contrast to the traditional MAML approach, which uses a gradient-based adaptation step, we propose employing a Hypernetwork to anticipate the values directly from the data.
The parameters for the final target model are calculated with the formula in Eq. 1.
$\theta^{\prime}=\theta+\Delta_{\theta}=\theta+H(\mathbf{E}_{\mathcal{S}},\,\mathbf{E}_{\theta}).$ (1)
Training
For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$. The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$. During training, we first sample the batch of sounds $\mathcal{B}$ from the given dataset $X$. Next, for each task data, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1). Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
$\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$ (2)
where $\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\lambda_{t}L_{t}(x,\hat{x})+\lambda_{f}L_{f}(x,\hat{x})$, with $L_{t}$ is L1 loss in temporal space, $L_{f}$ is multi-resolution mel-scale STFT loss in frequency space (Yamamoto et al., 2020) and $\lambda$s are coefficients controlling influence of particular losses. The Enconder parameters, universal weight encoder, Hypernetwork, and global parameters $\theta$ represent the meta parameters of the system and are updated with stochastic gradient descent (SGD) by optimizing $\mathcal{L}_{HyperMAML}(f_{\theta})$.
4 Experiments
In this section, we perform an experimental analysis. We divide it into two subsections: analysis of the performance of different INR architectures and experimental results of our hypernetwork approach, FewSound.
Multiple metrics are used to evaluate the quality of reconstructions. That includes quantitative metrics such as MSE, Log-spectral distance (LSD), PSNR, and SI-SNR. We also use metrics that measure the perceptual quality of the audio signal, such as PESQ (Rix et al., 2001), STOI (Taal et al., 2010), and CDPAM (Manocha et al., 2021). Given that the perception of ’natural’ sound is critical to our output, we prioritized the PESQ metric due to its strong correlation with human auditory perception.
4.1 Implicit Neural Representations
We compare the aforementioned INR architectures using two subsets of the VCTK (Veaux et al., 2013) dataset. Each of the subsets is generated by randomly choosing one recording per speaker, resampling it to 22.05 kHz and randomly cropping it to a desired length. We crop the recordings to 32,768 samples (about 1.5 seconds) in the first subset and to 65,536 (about 3 seconds) in the second subset. Both subsets consist of a total of 110 recordings. To make a fair comparison, all of the MLP based INRs have the same number of layers and the same sizes of layers, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models. All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov & Hutter, 2019) and the same loss function as in FewSound. Table 1 shows the results of the experiments. All INRs were selected to be close in the number of parameters. One may note that, with the exception of MSE, the shorter audios result in better quality. KAN is competitive for most metrics, achieving the best mean score in LSD, PESQ, STOI and CDPAM and second best PSNR for 1.5 s audio; as well as the best mean LSD, PESQ, STOI, and second best MSE, PSNR, SI-SNR, CDPAM for 3 s audio.
4.2 Hypernetworks
Setting
When using hypernetworks, we only consider NeRF, SIREN, and KAN as the target network architectures. We compare our method with the state-of-the-art model HyperSound (Szatkowski et al., 2023) and follow the same experimental setting. We use the VCTK dataset resampled to $22,050$ Hz. The inputs have a length of 32,768 samples and are obtained by cropping the original recordings. The last 10 speakers are used as the validation set. All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch. We use AdamW optimizer (Loshchilov & Hutter, 2019) and OneCycleLR scheduler (Smith & Topin, 2018). The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is set to $1\mathrm{e}{-6}$.
We examine how different encoder and target network architectures impact the quality of reconstructions. In particular, we explore the performance of KAN-based target networks used both with FewSound and HyperSound models.
Unless otherwise mentioned, FewSound uses a SoundStream (Zeghidour et al., 2021) based encoder, 17.5M parameter hypernetwork MLP and NERF-based target network. The number of parameters in target networks is roughly equal to the length of the input audio (32K). Results for HyperSound are taken from the original paper.
Results
Table 2 shows the comparison of our model with state-of-the-art HyperSound. We include results for two sizes of the hypernetwork MLP – a smaller 17.5M parameter model and a larger one with 41.2M parameters. Both of them have comparable results, showing that the smaller network may be used for efficiency. One may note that even the smaller one outperforms HyperSound in all the metrics (MSE by 33.3%, LSD by 6.83%, SI-SINR by 60.87%, PESQ by 8.66%, STOI by 2.63% and CDPAM by 13.89%).
Table 2: FewSound performance compared to HyperSound.
Model MSE LSD SI-SNR PESQ STOI CDPAM
$\rightarrow 0$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
HyperSound 0.009 1.61 2.99 1.27 0.76 0.36
FewSound 17.5M 0.006 1.50 4.81 1.38 0.78 0.31
FewSound 41.2M 0.006 1.50 5.03 1.41 0.79 0.30
Impact of target network architecture
Tables 3 and 4 show how different target network architectures impact the quality of reconstructions using FewSound and HyperSound models, respectively. Note that for FewSound NERF outperforms other architectures in terms of all metrics, with KAN close second (between 1.1% and 16% worse). Whereas, for HyperSound, KAN is better in 4 out of 6 metrics (between 1.3% and 28.5%) and worse in two (0.6% and 11%). For all the experiments, SIREN performs the worst.
Table 3: Performance of FewSound with different target network architectures.
MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
NERF 0.006 28.07 1.50 4.81 1.38 0.78 0.31
KAN 0.007 27.76 1.55 4.26 1.36 0.77 0.36
SIREN 0.025 22.02 1.74 -9.40 1.18 0.64 0.31
Table 4: Performance of HyperSound with different target network architectures. Results for SIREN and NERF are taken from the original paper, while we train the model using the KAN target network architecture ourselves. A target network with a significantly lower number of parameters (19K) is used in this case, as it achieves better results.
MSE LSD SI-SNR PESQ STOI CDPAM
$\rightarrow 0$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
NERF 0.009 1.61 2.99 1.27 0.76 0.36
KAN 0.007 1.62 3.98 1.29 0.77 0.40
SIREN 0.018 3.56 -3.66 1.19 0.65 0.3
Ablation Study
Table 5 shows the impact of using different encoder architectures, mentioned in Section 2. SoundStream, SNAC, and Spectrogram Encoder were trained from scratch. In the case of Whisper, we base on the pre-trained checkpoints - tiny.en and large-v2 and finetune using the LoRA approach (Hu et al., 2022).
Table 5: FewSound performance with different encoder architectures.
Encoder MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
SoundStream 0.006 28.07 1.50 4.81 1.38 0.78 0.31
Whisper L 0.042 19.72 1.40 -27.67 1.30 0.75 0.31
Whisper T 0.039 20.10 1.40 -29.08 1.25 0.72 0.33
Spectral 0.045 19.46 1.35 -21.76 1.49 0.80 0.29
SNAC 0.007 27.55 1.56 4.03 1.30 0.77 0.32
Table 6: KAN performance on the 1.5 s samples, with respect to the encoding length. The results are averaged across 10 samples.
Enc. Len. #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
2 23,016 0.0033 30.7170 2.3927 9.0285 1.4980 0.7556 0.3731
4 25,704 0.0017 33.7562 1.6014 12.3093 1.9463 0.9040 0.2561
8 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
10 33,768 0.0005 41.1152 1.3921 20.0078 3.4947 0.9930 0.1574
12 36,456 0.0005 42.6504 1.4182 21.6062 3.5176 0.9949 0.1664
16 41,832 0.0002 46.8576 1.4194 25.9786 3.6256 0.9957 0.1538
24 52,584 0.0003 45.7863 1.4764 24.9128 3.6033 0.9955 0.1982
Table 7: KAN results on 1.5 s samples in relation to different number and sizes of hidden layers.
Layer Sizes #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
[24, 12, 6] 10,500 0.0014 35.9192 1.4806 14.5340 2.1313 0.9360 0.2977
[24, 24, 12, 6] 18,564 0.0012 37.6839 1.3589 16.4163 2.7035 0.9778 0.2746
[48, 12, 6] 19,908 0.0010 38.5736 1.3589 17.3090 2.8569 0.9795 0.2679
[48, 12, 12, 6] 21,924 0.0010 38.2946 1.3557 17.0775 2.9597 0.9810 0.2618
[48, 16, 8] 23,408 0.0009 38.7324 1.3347 17.5094 3.0476 0.9845 0.2451
[48, 24, 12] 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
[96, 48, 24] 102,480 0.0003 43.6428 1.1715 22.5679 4.0203 0.9964 0.1072
We further analyze the influence of various parameters’ values on the KAN’s performance. In the ablation study, experiments are also conducted on audio samples of 1.5 seconds and 3 seconds, with each subset limited to only 10 recordings. Due to the page limit, we include selected results for 1.5 s audio. More parameters and analysis of 3 s audio samples can be found in Appendix A. During our study we keep the following base parameters’ values, changing only the analyzed components: encoding length = 8, grid size = 10, spline order = 2, layer sizes = [48, 24, 12], scale noise = 0.1, scale base = 1, scale spline = 1, learning rate = 5e-3 and enable scale spline = True, which controls whether we include the learnable spline weigths matrix $w_{s}$. Theoretically, these weigths are redundant, as they could be absorbed into the spline itself.
Table 6 illustrates the impact of the length of the encoding (size of the positional encoding) that directly influences the size of networks. Values smaller than the baseline of 8 result in progressively worse performance; meanwhile, increasing the length provides improvement up to value of 16, where the results are the best for all but one metric (LSD).
Table 7 demonstrates the impact of different hidden layer configurations. The results indicate a consistent performance improvement as network capacity increases, with the architecture containing hidden layers of [96, 48, 24] achieving the best performance. Decreasing the networks size leads to progressive results degradation.
Table 8 presents a comparison of the varying spline order values (the degree of the applied spline). Increasing the spline order allows creation of more complex spline functions; however, this comes at the cost of additional parameters. In our study, an order of 3 provides the best overall performance across both quantitative (PSNR) and perceptual (PESQ) metrics. Using smaller or larger spline orders leads to a decline in metric scores.
Table 9 provides insights into the influence of grid size (number of knots defining the spline) on KAN’s learning process. Similarly to encoding length and spline order, an increase in grid size leads to a higher number of INR parameters. While a larger spline order eventually results in a decline in reconstruction quality, this effect is not observed with grid size. Extra parameters enhance the expressiveness of the INR, improving all metrics except LSD.
Table 8: KAN spline order (O) for 1.5 s audio recordings
O Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
1 28,860 9e-4 38.83 1.53 17.59 3.01 0.98 0.27
2 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
3 33,300 7e-4 40.43 1.23 19.28 3.46 0.99 0.18
5 37,740 1e-3 38.87 1.22 17.66 3.38 0.99 0.19
7 42,180 9e-4 38.91 1.23 17.71 3.21 0.98 0.21
Table 9: Impact of spline function grid size (G) on 1.5 s reconstructions
G Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
1 11,100 3e-3 30.65 2.00 8.99 1.65 0.80 0.43
5 19,980 1e-3 37.33 1.36 16.01 2.61 0.97 0.27
10 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
12 35,520 5e-4 42.34 1.30 21.24 3.65 0.99 0.16
17 46,620 3e-4 44.78 1.31 23.78 3.81 1.00 0.13
Spectrograms
The spectrograms shown in Fig. 2 illustrate differences in the fidelity of sound reconstruction achieved by various INRs compared to the ground truth. Across all methods, the harmonic structures in the low-frequency range are generally well-preserved, reflecting their ability to capture fundamental features of the audio signal. However, significant variations emerge in the representation of high-frequency components, where most INRs lowers the intensity of the components. It is particularly visible with FINER. Moreover, all of the INRs (with KAN and NERF to the least degree) demonstrate a tendency to over-smooth the higher intensities in high-frequency components, which results in adding white noise. Finally, one can note the lower average intensity in the reconstructed audio.
Figure 2: The comparison of spectrograms. We compare the original sample with the waveforms reconstructed using different INR approaches.
5 Conclusion
This work demonstrates the effectiveness of Kolmogorov-Arnold Networks (KANs) as implicit neural representations for audio signals, offering high reconstruction accuracy when integrated with hypernetwork-based architectures. By evaluating KANs against classical INR models like NeRF, FINER, WIRE, RFF, and SIREN, we highlight their potential as a foundational approach for efficient audio modeling, resulting in top-2 mean in 5 out of 7 considered quality metrics for 1.5 s audio, and top-2 in all 7 for 3 s audio. Additionally, by integrating KAN into hypernetwork frameworks like FewSound, performance improves significantly, surpassing HyperSound with a 33.3% reduction in MSE, a 60.87% increase in SI-SNR, and an 8.66% in PESQ. The results suggest that KANs can serve as a powerful tool for audio representation, enabling adaptive processing across various datasets and tasks.
Since our experiments were conducted at a fixed sampling rate of 22,050 Hz, the impact of varying sampling strategies requires investigation in future work. Furthermore, integration with modern neural audio codecs and speech encoder systems presented in Tab 5 might be further expanded, in particular when investigating multilingual approach. The evaluation was restricted to English datasets due their prevalence. Additionally, the datasets lacked complex auditory scenarios, such as multi-speaker dialogues, emotional speech, and diverse acoustic scenes, which might be an interesting extensions to the research. It is important to note, that while KANs demonstrated efficiency in certain tasks, their computational demands may grow significantly for longer sequences or large-scale applications.
Future research should extend KAN-based models to multilingual datasets and diverse linguistic settings to enhance their generalization. Investigating the impact of sample length on performance could provide deeper insights into the balance between computational efficiency and fidelity. Architectural refinements may reduce resource requirements while maintaining accuracy, making KANs more scalable for practical deployment. Exploring their integration with multimodal frameworks, such as facial expressions or gestures, could enable advanced audio-visual synthesis and applications like avatar generation, an interesting path given hypernetworks already being used in visual modality (Przewiezlikowski et al., 2024; Kania et al., 2024). Additionally, developing interpretable models would improve transparency and understanding of KANs’ decision-making processes, which would increase trust in real-world applications and meeting legal requirements for practical usage. Finally, optimizing compatibility with a broader range of audio encoding systems could significantly enhance reconstruction quality, expanding the practical utility of KAN-based audio representations.
Appendix A Extended ablation study
At the beginning of this section we conclude the ablation study for 1.5 s audio samples by analyzing the influence of learning rate, standalone scale spline flag, scale base factor, scale noise factor and scale spline factor on KAN’s performance.
Table 10: Effect of various learning rate values on 1.5 s recordings
Learning rate Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.01 31,080 0.0007 40.5596 1.3698 19.4728 3.2799 0.9877 0.1919
0.007 31,080 0.0007 40.5170 1.3284 19.3607 3.3883 0.9908 0.1927
0.006 31,080 0.0007 40.0159 1.3051 18.9209 3.3532 0.9911 0.1770
0.005 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
0.001 31,080 0.0011 37.8920 1.3626 16.6339 2.5975 0.9711 0.2314
Table 10 shows the comparison of learning rates. Generally, KAN performs well and stably with higher learning rate setting. A learning rate of 0.001 yields the worst scores across all entries in the table. A learning rate of 0.01 achieves the best results in the PSNR and Si-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores.
Table 11 shows the comparison of performance of KAN depending on whether the enable scale spline flag is set to False or True. The former results in skipping the theoretically redundant weigths. However, it leads to a drop in all metric scores except MSE, with perceptual metrics being affected the most.
Table 11: Enable scale spline flag ablation for 1.5 s audio source
Enable Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
False 28,860 0.0007 39.4067 1.5253 18.2321 2.7388 0.9653 0.2771
True 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
At the end of this paragraph, we consider hyperparameters related to KAN initialization. The weight matrices $w_{b}$ and $w_{s}$ are initialized according to Kaiming uniform schema - the vaules are drawn from $\mathcal{U}(-k,k)$ distrubution, where
$k=\sqrt{\frac{6}{(1+5a^{2})\cdot fan\_in}}$
and $a$ is equal to scale base or scale spline for $w_{b}$ and $w_{s}$ respectively. The $c_{i}$s parameters are initialized by sampling a noise from $\mathcal{N}(0,\sigma^{2})$ distribution, and solving an instance of the least squares problem, so that $spline(x)$ for a certain $x$ is as close to the sampled noise as possible. The scale noise parameter is the standard deviation $\sigma$ of the noise distribution.
Tables 12 and 13 show the ablation results for scale spline and scale base parameters respectively. In case of the first one, 0.5 seems to be the most optimal choice with the highest score for all metrics excluding LSD, where optimal choice is 1, and STOI, where the difference between the best scores is negligible. In case of scale spline the choice of the best value varies between metrics. Table 14 shows the results for the study of scale noise hyperparameter, with the lowest tested value of 0.025 attaining the best results in most of the metrics.
Table 12: KAN performance on 1.5 s, with respect to the scale base values.
Scale Base Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.25 31,080 0.0008 39.9112 1.3386 18.7366 3.3149 0.9888 0.1872
0.5 31,080 0.0006 41.0193 1.3111 19.8693 3.4322 0.9904 0.1875
1 31,080 0.0006 40.8162 1.2895 19.6861 3.3821 0.9900 0.1930
2 31,080 0.0006 40.5262 1.3129 19.4204 3.2141 0.9906 0.2013
4 31,080 0.0008 39.5773 1.3195 18.4513 3.2608 0.9876 0.2062
Table 13: KAN performance on 1.5 s, with respect to the scale spline values.
Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.25 31,080 0.0006 40.6718 1.3196 19.5770 3.5030 0.9917 0.1932
0.5 31,080 0.0005 40.9010 1.3057 19.8454 3.4086 0.9915 0.1895
1 31,080 0.0007 40.5384 1.2895 19.4369 3.4281 0.9915 0.1872
2 31,080 0.0006 41.5030 1.2973 20.3684 3.3479 0.9914 0.1815
4 31,080 0.0006 40.8259 1.2981 19.7534 3.3932 0.9903 0.1887
Table 14: KAN performance on 1.5 s, with respect to the scale noise values.
KAN Scale Noise Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.4 31,080 0.0007 40.9740 1.3010 19.8608 3.4433 0.9910 0.1896
0.2 31,080 0.0007 40.2776 1.2849 19.1855 3.4280 0.9915 0.1796
0.1 31,080 0.0007 40.5541 1.2796 19.4436 3.4576 0.9913 0.1897
0.05 31,080 0.0007 40.0844 1.2864 18.9432 3.4004 0.9909 0.1846
0.025 31,080 0.0006 41.0045 1.2878 19.9189 3.4605 0.9897 0.1825
Performance analysis of KAN on 3 s sounds
Next, we present the ablation study for 3 s recordings, considering the same hyperparameters as used for the 1.5 s samples.
Table 15 shows the impact of different encoding lengths on 3 s utterances. The absolute performance is generally lower across all metrics in relation to 1.5 s samples, e.g. the best PESQ decreasing from 3.62 to 2.90 and the best SI-SNR decreasing from 25.97 to 17.17. The results are much less conclusive than in case of 1.5 s, where length 16 was clearly optimal across most metrics. However, the majority of metrics (MSE, PSNR, SI-SNR and PESQ, STOI) are the best for encoding lengths of 24 and 16 respectively.
Table 15: KAN performance on the 3 s samples, with respect to the encoding length.
Encoding Length Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
2 23,016 0.0036 30.4569 3.3098 4.2598 1.2121 0.6627 0.4217
4 25,704 0.0011 35.4080 2.0309 10.3129 1.5506 0.8204 0.2926
8 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
10 33,768 0.0006 39.3990 1.3152 14.5468 2.8272 0.9779 0.3795
12 36,456 0.0005 39.6602 1.3722 14.7926 2.7818 0.9799 0.3730
16 41,832 0.0004 40.7796 1.4383 15.9502 2.9095 0.9820 0.3516
24 52,584 0.0003 41.9264 1.6033 17.1797 2.8385 0.9816 0.3429
Table 16 shows how sizes and depths of hidden layers influence results on 3 s samples. Similarly to 1.5 s (Table 7) the best performance is achieved by the largest network composed of [96, 48, 24] layers. The performance, however does not improve with the increasing size as gradually as for 1.5 s samples.
Table 16: KAN results on 3 s samples in relation to different number and sizes of hidden layers.
Layer Sizes Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
[24, 12, 6] 10,500 0.0007 37.5744 1.4904 12.6576 1.7192 0.8912 0.3229
[24, 24, 12, 6] 18,564 0.0006 38.4196 1.3828 13.5231 1.9078 0.9393 0.3847
[48, 12, 6] 19,908 0.0008 38.0106 1.3325 13.0259 2.1609 0.9384 0.3734
[48, 12, 12, 6] 21,924 0.0006 38.3097 1.3430 13.3983 2.1817 0.9517 0.3886
[48, 16, 8] 23,408 0.0007 38.2895 1.3119 13.3530 2.2985 0.9562 0.3756
[48, 24, 12] 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
[96, 48, 24] 102,480 0.0005 40.9599 1.1792 16.1224 3.5224 0.9898 0.1594
In comparison to the 1.5 s case, where a spline order of size 3 is clearly the optimal hyperparameter setting, its advantage is less evident in Table 17. KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI columns.
Table 17: KAN effectiveness in terms of diverse spline order sizes and 3 s long input
Spline Order Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
1 28,860 0.0006 38.5038 1.4711 13.6106 2.2903 0.9540 0.3299
2 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
3 33,300 0.0008 38.6639 1.2540 13.6820 2.5639 0.9704 0.3578
5 37,740 0.0006 39.0929 1.2663 14.1943 2.4499 0.9661 0.3376
7 42,180 0.0006 38.8824 1.3175 13.9657 2.1987 0.9573 0.3378
In table 18, table 19 and table 20, which correspond respectively to the ablations of grid size, learning rate, and enable scale spline flag on 3 s speeches, we observe similar conclusions and tendencies as in the 1.5s scenario. Described in Table 9 (Sect. 4.2), and Tables 10, 11(Appendix A), respectively.
Table 18: KAN effectiveness in terms of diverse grid size and 3 s long input
Grid size Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
1 11,100 0.0029 31.3359 2.0005 5.4555 1.3828 0.7652 0.3998
5 19,980 0.0007 38.1135 1.4668 13.2103 1.9651 0.9160 0.2399
10 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
12 35,520 0.0005 39.4903 1.2745 14.6419 2.6706 0.9739 0.3204
17 46,620 0.0006 39.7273 1.3224 14.8089 3.0683 0.9734 0.2646
Table 19: Effect of various learning rate values on 3 s recordings
Learning rate Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.01 31,080 0.0006 39.4961 1.3238 14.6358 2.5597 0.9692 0.3940
0.007 31,080 0.0006 39.2295 1.2804 14.3368 2.5669 0.9707 0.3685
0.006 31,080 0.0006 39.1037 1.2781 14.2218 2.5516 0.9700 0.3633
0.005 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
0.001 31,080 0.0007 37.7570 1.4674 12.8016 1.9302 0.9261 0.3020
Table 20: Enable scale spline flag ablation for 3 s audio source
Enable Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
False 28,860 0.0007 38.0986 1.4767 13.2310 2.1969 0.9360 0.4218
True 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
Finally, in Tables 21, 22, and 23 one can notice that the scale base and the scale spline hyperparameters do not significantly influence the performance. While manipulating scale base, all metrics except for CDPAM vary by less than 9% (the mentioned CDPAM under 16%). Moreover, apart from MSE that almost does not change, the changes in metric scores are not monotonic with respect to the changes in parameter values.
Table 21: Evaluation of different scale base factors on 3 s audio samples
Scale Base Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.25 31,080 0.0006 39.0743 1.2804 14.1621 2.4855 0.9654 0.3574
0.5 31,080 0.0006 39.1813 1.2744 14.2891 2.5813 0.9714 0.3485
1 31,080 0.0006 38.9647 1.2643 14.0714 2.5415 0.9693 0.3573
2 31,080 0.0006 38.8621 1.2980 13.9575 2.5120 0.9654 0.3932
4 31,080 0.0005 39.7455 1.3119 14.9256 2.3783 0.9612 0.4040
Table 22: Evaluation of different scale spline factors on 3 s audio samples
Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.25 31,080 0.0006 39.3700 1.2761 14.4961 2.6388 0.9713 0.3758
0.5 31,080 0.0006 38.9780 1.2698 14.0723 2.5493 0.9689 0.3678
1 31,080 0.0006 39.0889 1.2654 14.2169 2.4618 0.9643 0.3667
2 31,080 0.0006 38.9006 1.2735 13.9963 2.5288 0.9696 0.3663
4 31,080 0.0006 38.9421 1.2740 14.0168 2.4376 0.9666 0.3448
Table 23: Evaluation of different scale noise factors on 3 s audio samples
Scale Noise Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.4 31,080 0.0006 39.0755 1.2712 14.1628 2.5252 0.9659 0.3711
0.2 31,080 0.0006 39.1540 1.2631 14.2794 2.5684 0.9687 0.3604
0.1 31,080 0.0006 39.0749 1.2624 14.1858 2.5053 0.9698 0.3488
0.05 31,080 0.0006 39.0345 1.2699 14.1197 2.5075 0.9680 0.3601
0.025 31,080 0.0006 39.1100 1.2706 14.2064 2.5701 0.9679 0.3613

### v2

@@
 A Hypernetwork-Based Approach to KAN Representation of Audio Signals
-Abstract
+Abstract.
 Implicit neural representations (INR) have gained prominence for efficiently encoding multimedia data, yet their applications in audio signals remain limited.
@@
 KAN demonstrates superior perceptual performance over previous INRs, achieving the lowest Log-Spectral Distance of 1.29 and the highest Perceptual Evaluation of Speech Quality of 3.57 for 1.5 s audio.
-To extend KAN’s utility, we propose FewSound [^1] The source code can be accessed at https://github.com/gmum/fewsound.git, a hypernetwork-based architecture that enhances INR parameter updates.
+To extend KAN’s utility, we propose FewSound, a hypernetwork-based architecture that enhances INR parameter updates.
 FewSound outperforms the state-of-the-art HyperSound, with a 33.3% improvement in MSE and 60.87% in SI-SNR.
@@
 representation learning, audio representation, INR, KAN
-1 Introduction
-Implicit neural representations (INRs) represent discrete data in neural network weights.
+conference: ; isbn: ;ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Learning paradigmsccs: Applied computing Sound and music computing
+Figure 1.
+Sound representation is a challenging task.
+In classical analysis, we usually apply the FFT (Fast Fourier Transform), which analyzes frequencies by separating them into sinusoidal components.
+We propose to analyze sounds using KAN (Kolmogorov-Arnold Network), which uses trainable activation functions instead of classical layers.
+Such an operation can be seen as approximating signals by a linear combination of basis functions, bearing a similarity to the FFT or wavelets.
+1. Introduction
+Implicit neural representations (INRs) depict discrete data in neural network weights.
 INR finds many applications in the case of image (Klocek et al., 2019), video (Chen et al., 2022), or 3D objects (Spurek et al., 2020).
 In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data.
-HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes on raw input sound and produces weight for INR-based sound representation.
+HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes raw input sound and produces weight for INR-based sound representation.
 INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes.
-Siamese SIREN (Lanzendörfer & Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
+Siamese SIREN (Lanzendörfer and Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
 The recently proposed KAN (Kolmogorov-Arnold Network) (Liu et al., 2024a) replaces traditional fully connected layers with activation functions.
@@
 However, its effectiveness in low-dimensional settings and successful application to complex function regression tasks make it a promising candidate for audio INR.
-This paper demonstrates that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm.
+This paper shows that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm.
 For single audio INR, KAN trains more rapidly and achieves better results.
@@
 In practice, we can directly produce KAN parameters or use updates for the universal KAN-based architecture.
-Therefore, we propose two methods.
-The first is a classical HyperSound (Szatkowski et al., 2023) with a KAN target network.
+Hence, we propose two methods, a classical HyperSound (Szatkowski et al., 2023) with a KAN target network.
 The second uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR, which is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017).
-Since in a few-shot approach, there does not exist a model dedicated to audio signals, we introduce FewSound.
-FewSound takes raw sound on input and uses hypernetwork to generate updates to universal weights of INR dedicated to encoding sound.
+As there is no few-shot model for audio signals, we introduce FewSound  [^1] The source code for the proposed method can be found at the following link: https://github.com/gmum/fewsound.
+It takes raw sound on input and uses a hypernetwork to generate updates to the universal weights of INR dedicated to encoding sound.
 In our work, we examine six approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a).
@@
 Thus, two settings show that KAN can be effectively modeled by hypernetworks.
-Our study demonstrates that KAN model can be viewed as INR of sound.
+Our study demonstrates that KAN can be viewed as INR of sound.
 Such models yield favorable outcomes when dealing with individual audio signals and excel in large Hypernetwork-based models trained on large datasets.
+We use KANs due to their spline-based activations.
+Since any function f (e.g., an audio representation of the magnitude over time) can be approximated by splines with $\sup\|f\_spline\|=\mathrm{O}(h^{4})$ (Hall and Meyer, 1976), KANs inherit this theoretical precision as a linear combination of splines, resulting in a polynomial bound.
+This suggests that KANs could be effective for audio reconstruction.
 Key contributions of our research include:
-  • showing KAN’s efficacy for continuous sound representation,
+  • introducing KAN-based sound representation and showing its efficacy,
+  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN network and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
   • comparing KAN represenation with other INRs,
   • showing that KAN can be produced via hypernetworks,
-  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN network and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
-  • analysis of the parameters’s influence on the efficacy of the proposed method.
-2 Related Work
+  • analysis of the parameters’ influence on the efficacy of the proposed method.
+2. Related Work
 To our knowledge, there are no specialized INR architectures focused on sound.
@@
 Such information is limited by the networks’ architecture instead of the grid resolution.
-INRs are used in many signal processing problems, including 3D reconstruction (Mildenhall et al., 2020), super-resolution (Chen et al., 2022), compression (Strümpler et al., 2022; Lanzendörfer & Wattenhofer, 2023) or audio scene representation (Su et al., 2022).
+INRs are used in many signal processing problems, including 3D reconstruction (Mildenhall et al., 2020), super-resolution (Chen et al., 2022), compression (Strümpler et al., 2022; Lanzendörfer and Wattenhofer, 2023) or audio scene representation (Su et al., 2022).
 Initially, the INRs used ReLU-based MLPs.
@@
 It uses a variable-periodic activation $a(x)=sin(\omega_{0}\lvert x+1\rvert x)$ and a specialized bias initialization schema in its first layer.
-NERF (Neural Radiance Field) (Mildenhall et al., 2020) is a type of network originally designed for scene representation.
-Based on the 5D input of spatial location $(x,\,y,\,z)$ and viewing direction ($\theta,\,\phi$), the MLP network outputs color c = $(r,\,g,\,b)$ and volume density ($\sigma$).
-This allows the generation of 3D reconstructions of scenes from a set of sparse views.
+NeRF (Neural Radiance Field) (Mildenhall et al., 2020) models scenes via a 5D input—spatial location $(x,\,y,\,z)$ and viewing direction $(\theta,\,\phi)$—with an MLP outputting color $c=(r,\,g,\,b)$ and density $\sigma$, enabling 3D reconstruction from sparse views.
+Figure 2.
+The overview of FewSound architecture.
+The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights.
+The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$.
+The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
 Sound representation
@@
 In our experiments we used only the encoder part composed of 1D convolution followed by four encoder blocks with residual units and the final 1D convolution.
-SNAC (Siuzdak, 2024) uses the encoder based on the hierarchical tokens like the previous approach.
-The distinction is that the coarse tokens are sampled less frequently, resulting in covering broader time windows.
-SNAC not only decreases the bitrate required for reconstruction, but also provides useful information in language modeling tasks.
+SNAC (Siuzdak et al., 2024) employs a hierarchical token-based encoder, similar to prior work, but samples coarse tokens less frequently to span longer time windows.
+This reduces reconstruction bitrate and aids language modeling tasks.
 Whisper (Radford et al., 2023) is a Transformer-based (Vaswani et al., 2017) ASR system, trained on 680,000 hours of data, achieving state-of-the-art speech recognition results.
@@
 We consider six variants of INRs: NeRF, SIREN, RFF, FINER, WIRE and KAN.
-We use NeRF to refer to a network with positional encoding, followed by a MLP with ReLU activation.
-The positional encoding, described by $\gamma(t)=\left[\sin(2^{0}\pi t),\cos(2^{0}\pi t),\ldots,\sin(2^{L-1}\pi t),\cos(2^{L-1}\pi t)\right]$ is taken from the original NeRF (Mildenhall et al., 2020).
-SIREN (Sitzmann et al., 2020) is a MLP with sine activation: $a(x)=\sin(\omega_{0}x)$, where $\omega_{0}$ is a hyperparameter.
-RFF (Random Fourier Features) (Tancik et al., 2020) is used to refer to a multilayer perceptron with ReLU activation and input embedding defined as $\gamma(t)=\left[\cos(2\pi\mathbf{B}t),\sin(2\pi\mathbf{B}t)\right],$ where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is a random projection matrix, sampled from $\mathcal{N}(0,\sigma^{2})$.
+We use NeRF to refer to a network with positional encoding, followed by an MLP with ReLU activation.
+The positional encoding, described by
+$\gamma(t)=\left[\sin(2^{0}\pi t),\cos(2^{0}\pi t),\ldots,\sin(2^{L-1}\pi t),\cos(2^{L-1}\pi t)\right]$
+is taken from the original NeRF (Mildenhall et al., 2020).
+SIREN (Sitzmann et al., 2020) is an MLP with sine activation: $a(x)=\sin(\omega_{0}x)$, where $\omega_{0}$ is a hyperparameter.
+RFF (Random Fourier Features) (Tancik et al., 2020) is used to refer to an MLP with ReLU activation and input embedding defined as
+$\gamma(t)=\left[\cos(2\pi\mathbf{B}t),\sin(2\pi\mathbf{B}t)\right],$
+where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is a random projection matrix, sampled from $\mathcal{N}(0,\sigma^{2})$.
 In our case $d$, the dimensionality of the input, is always equal to 1, while $\sigma$ and $m$ are hyperparameters.
-WIRE (Saragadam et al., 2023) is a MLP with complex Gabor wavelet activation $a(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$, where $\omega_{0}$ and $s_{0}$ are hyperparameters.
-FINER (Liu et al., 2024b) is a MLP with activation function defined as $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$, where $\omega_{0}$ is a hyperparameter.
+WIRE (Saragadam et al., 2023) is an MLP with complex Gabor wavelet activation $a(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$, where $\omega_{0}$ and $s_{0}$ are hyperparameters.
+FINER (Liu et al., 2024b) is an MLP with activation function defined as $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$, where $\omega_{0}$ is a hyperparameter.
 The bias of its first layer is initialized by sampling from $\mathcal{U}(-k,k)$, where $k$ is a hyperparameter.
 Our Kolmogorov-Arnold network implementation consists of the same positional encoding as in NeRF, followed by KAN layers.
-Each KAN layer can be described by $\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}spline(x),$ where $w_{b}$ and $w_{s}$ are matrices of learnable weights, and $spline(x)$ is a linear combination of B-splines such that $spline(x)=\sum_{i}{c_{i}B_{i}(x)},$ with learnable $c_{i}$s.
-It is parametrized by spline order – the degree of the used spline, and grid size – the number of knots defining the spline.
-3 KAN in hypernetwork-based architectures
+Each KAN layer can be described by $\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}spline(x),$ where $w_{b}$ and $w_{s}$ are matrices of learnable weights, and $spline(x)$ is a linear combination of B-splines such that $spline(x)=\sum_{i}{c_{i}B_{i}(x)}$ with learnable $c_{i}$s.
+It is parametrized by spline order – the degree of the spline, and grid size – the number of knots defining the spline.
+3. KAN in hypernetwork-based architectures
 In this section, we show how KAN works with hypernetworks.
@@
 Drawing a parallel between hypernetworks and generative models, the authors of (Sheikh et al., 2017) employ this technique to generate a variety of target networks that approximate the same function.
-Figure 1: The overview of FewSound architecture.
-The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights.
-The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$.
-The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
-Table 1: Reconstruction metrics for 1.5 s and 3 s audio. We report mean and standard deviation across 110 samples, each of a different speaker of VCTK dataset (Veaux et al., 2013).
+Few-shot learning for INR
+A significant limitation of deep learning models is their reliance on extensive datasets.
+In contrast, humans can learn effectively from just a few examples.
+A few-shot learning (Wang et al., 2020) approach was proposed to bridge the gap between deep neural network training and human learning capabilities.
+This strategy treats the data set as consisting of numerous small tasks.
+Few-shot architectures can be quickly adapted to solve new problems in a single step.
+The most popular approach to few-shot learning is MAML (Finn et al., 2017).
+MAML uses a set of universal weights, fine-tuned for specific small tasks through a limited number of gradient updates.
+Due to the model-agnostic nature of the approach, it can be adapted to various models and corresponding learning algorithms.
+The terminology of few-shot learning is often inconsistent due to varying definitions in existing research.
+For a standardized framework, see  (Chen et al., 2019; Wang et al., 2020).
+However, such a nomenclature is dedicated to a classification, regression, or generative model task.
+In the case of the INR-based model, we refer to (Tancik et al., 2021).
+We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
+FewSound
+Now we are ready to present FewSound – a model that uses Hypernetworks to update the INR weights dedicated to sound.
+The core concept of this update mechanism is to use information extracted from raw sound data and universal weights to determine the optimal updates for a particular task.
+This method adjusts the INR’s parameters for entirely different sounds.
+The schema of FewSound is provided in Fig. 2. We aim to reconstruct the training example using the target network. In practice, we run our INR on all steps to reconstruct the input.
+Following our previous analysis of INR in the sound reconstruction task, we consider the parameterized function $f_{\theta}$ to be one of the discussed INRs.
+In addition, we distinguish the trainable encoding network $E(\cdot)$ in our architecture, which transforms raw audio data into low-dimensional representation.
+We investigated various encoder architectures and report their performance in Tab.
+5.
+The encoded sound is concatenated with universal weights propagated by the universal weight encoder $\mathbf{G}(\cdot)$ and then to the fully connected hypernetwork $H(\cdot)$.
+The hypernetwork then modifies the universal weights: $\theta^{\prime}=\theta+\Delta_{\theta}.$ In contrast to the traditional MAML approach, which uses a gradient-based adaptation step, we propose employing a Hypernetwork to anticipate the values directly from the data.
+Target model parameters are computed via Eq. 1.
+(1) $\theta^{\prime}=\theta+\Delta_{\theta}=\theta+H(\mathbf{E}_{\mathcal{S}},\,\mathbf{E}_{\theta}).$
+While FewSound accepts a fixed-size input (in our case 32768), it can be easily used on audio samples of arbitrary length by reconstructing them piece by piece and smoothing during merging.
+Training
+For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$.
+The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$.
+During training, we first sample the batch of sounds $\mathcal{B}$ from the dataset $X$.
+Next, for each task, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1).
+Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
+(2) $\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$
+where $\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\lambda_{t}L_{t}(x,\hat{x})+\lambda_{f}L_{f}(x,\hat{x})$, with $L_{t}$ is L1 loss in temporal space, $L_{f}$ is multi-resolution mel-scale STFT loss in frequency space (Yamamoto et al., 2020) and $\lambda$s are coefficients controlling influence of particular losses.
+The Enconder parameters, universal weight encoder, Hypernetwork, and global parameters $\theta$ represent the meta parameters of the system and are updated with stochastic gradient descent (SGD) by optimizing $\mathcal{L}_{HyperMAML}(f_{\theta})$.
+Table 1. Reconstruction metrics for 1.5 s and 3 s audio. We report mean and standard deviation across 110 samples, each of a different speaker of VCTK dataset (Veaux et al., 2013).
 1.5 s audio
@@
 KAN (33,768) 3.8e-4 $\pm$ 4.4e-4 42.28 $\pm$ 5.31 1.34 $\pm$ 0.19 18.85 $\pm$ 5.08 2.89 $\pm$ 0.51 0.97 $\pm$ 0.03 0.33 $\pm$ 0.15
-Few-shot learning for INR
-A significant limitation of deep learning models is their reliance on extensive datasets.
-In contrast, humans can learn effectively from just a few examples.
-A few-shot learning (Wang et al., 2020) approach was proposed to bridge the gap between deep neural network training and human learning capabilities.
-This strategy treats the data set as consisting of numerous small tasks.
-Few-shot architectures can be quickly adapted to solve new problems in a single step.
-The most popular approach to few-shot learning is MAML (Finn et al., 2017).
-MAML uses a set of universal weights, fine-tuned for specific small tasks through a limited number of gradient updates.
-Due to the model-agnostic nature of the approach, it can be adapted to various models and corresponding learning algorithms.
-The terminology of few-shot learning is often inconsistent due to varying definitions in existing research.
-For a standardized framework (Chen et al., 2019; Wang et al., 2020).
-However, such a nomenclature is dedicated to a classification, regression, or generative model task.
-In the case of the INR-based model, we refer to (Tancik et al., 2021).
-We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
-FewSound
-Now we are ready to present FewSound – a model that uses Hypernetworks to update the INR weights dedicated to sound.
-The core concept of this update mechanism is to use information extracted from raw sound data and universal weights to determine the optimal updates for a particular task.
-This method adjusts the INR’s parameters for entirely different sounds.
-The schema of FewSound is provided in Fig. 1. We aim to reconstruct the training example using the target network. In practice, we run our INR on all steps to reconstruct the input.
-Following our previous analysis of INR in the sound reconstruction task, we consider the parameterized function $f_{\theta}$ to be one of the discussed INRs.
-In addition, we distinguish the trainable encoding network $E(\cdot)$ in our architecture, which transforms raw audio data into low-dimensional representation.
-We investigated various encoder architectures and report their performance in Tab.
-5.
-The encoded sound is concatenated with universal weights propagated by the universal weight encoder $\mathbf{G}(\cdot)$ and then to the fully connected hypernetwork $H(\cdot)$.
-The hypernetwork then modifies the universal weights: $\theta^{\prime}=\theta+\Delta_{\theta}.$ In contrast to the traditional MAML approach, which uses a gradient-based adaptation step, we propose employing a Hypernetwork to anticipate the values directly from the data.
-The parameters for the final target model are calculated with the formula in Eq. 1.
-$\theta^{\prime}=\theta+\Delta_{\theta}=\theta+H(\mathbf{E}_{\mathcal{S}},\,\mathbf{E}_{\theta}).$ (1)
-Training
-For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$.
-The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$.
-During training, we first sample the batch of sounds $\mathcal{B}$ from the given dataset $X$.
-Next, for each task data, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1).
-Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
-$\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$ (2)
-where $\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\lambda_{t}L_{t}(x,\hat{x})+\lambda_{f}L_{f}(x,\hat{x})$, with $L_{t}$ is L1 loss in temporal space, $L_{f}$ is multi-resolution mel-scale STFT loss in frequency space (Yamamoto et al., 2020) and $\lambda$s are coefficients controlling influence of particular losses.
-The Enconder parameters, universal weight encoder, Hypernetwork, and global parameters $\theta$ represent the meta parameters of the system and are updated with stochastic gradient descent (SGD) by optimizing $\mathcal{L}_{HyperMAML}(f_{\theta})$.
-4 Experiments
-In this section, we perform an experimental analysis.
-We divide it into two subsections: analysis of the performance of different INR architectures and experimental results of our hypernetwork approach, FewSound.
+4. Experiments
+This section presents an experimental analysis in two parts: performance comparison of INR architectures and results of our hypernetwork method, FewSound.
 Multiple metrics are used to evaluate the quality of reconstructions.
@@
 Given that the perception of ’natural’ sound is critical to our output, we prioritized the PESQ metric due to its strong correlation with human auditory perception.
-4.1 Implicit Neural Representations
+4.1. Implicit Neural Representations
 We compare the aforementioned INR architectures using two subsets of the VCTK (Veaux et al., 2013) dataset.
@@
 Both subsets consist of a total of 110 recordings.
-To make a fair comparison, all of the MLP based INRs have the same number of layers and the same sizes of layers, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models.
-All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov & Hutter, 2019) and the same loss function as in FewSound.
+To make a fair comparison, all of the MLP based INRs have the same number of layers and with the same sizes, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models.
+All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov and Hutter, 2019) and the same loss function as in FewSound.
 Table 1 shows the results of the experiments.
-All INRs were selected to be close in the number of parameters.
 One may note that, with the exception of MSE, the shorter audios result in better quality.
 KAN is competitive for most metrics, achieving the best mean score in LSD, PESQ, STOI and CDPAM and second best PSNR for 1.5 s audio; as well as the best mean LSD, PESQ, STOI, and second best MSE, PSNR, SI-SNR, CDPAM for 3 s audio.
-4.2 Hypernetworks
+Further comparison of these INRs on additional datasets is available in section A.2.
+4.2. Hypernetworks
 Setting
@@
 We use the VCTK dataset resampled to $22,050$ Hz.
-The inputs have a length of 32,768 samples and are obtained by cropping the original recordings.
+The inputs are truncated to length of 32,768 samples.
 The last 10 speakers are used as the validation set.
 All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch.
-We use AdamW optimizer (Loshchilov & Hutter, 2019) and OneCycleLR scheduler (Smith & Topin, 2018).
-The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is set to $1\mathrm{e}{-6}$.
+We use AdamW optimizer (Loshchilov and Hutter, 2019) and OneCycleLR scheduler (Smith and Topin, 2018).
+The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is $1\mathrm{e}{-6}$.
 We examine how different encoder and target network architectures impact the quality of reconstructions.
@@
 One may note that even the smaller one outperforms HyperSound in all the metrics (MSE by 33.3%, LSD by 6.83%, SI-SINR by 60.87%, PESQ by 8.66%, STOI by 2.63% and CDPAM by 13.89%).
-Table 2: FewSound performance compared to HyperSound.
+Table 2. FewSound performance compared to HyperSound.
 Model MSE LSD SI-SNR PESQ STOI CDPAM
@@
 For all the experiments, SIREN performs the worst.
-Table 3: Performance of FewSound with different target network architectures.
-MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+Table 3. Effectiveness of FewSound with different target network architectures.
+Target MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
@@
 SIREN 0.025 22.02 1.74 -9.40 1.18 0.64 0.31
-Table 4: Performance of HyperSound with different target network architectures.
+Table 4.
+Performance of HyperSound with different target network architectures.
 Results for SIREN and NERF are taken from the original paper, while we train the model using the KAN target network architecture ourselves.
 A target network with a significantly lower number of parameters (19K) is used in this case, as it achieves better results.
-MSE LSD SI-SNR PESQ STOI CDPAM
-$\rightarrow 0$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
+Target MSE LSD SI-SNR PESQ STOI CDPAM
+Network $\rightarrow 0$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 NERF 0.009 1.61 2.99 1.27 0.76 0.36
@@
 SIREN 0.018 3.56 -3.66 1.19 0.65 0.3
-Ablation Study
+Influence of encoder architecture
 Table 5 shows the impact of using different encoder architectures, mentioned in Section 2.
-SoundStream, SNAC, and Spectrogram Encoder were trained from scratch.
+SoundStream, SNAC, and Spectrogram Encoder are trained from scratch.
 In the case of Whisper, we base on the pre-trained checkpoints - tiny.en and large-v2 and finetune using the LoRA approach (Hu et al., 2022).
-Table 5: FewSound performance with different encoder architectures.
+Among the encoders considered, SoundStream achieves the best results in terms of distance-based metrics such as PSNR, while Spectrogram Encoder excels in perceptual ones.
+SNAC doesn’t yield poor results, but it is essentially a less effective version of SoundStream.
+The Whisper models, by contrast, show significantly lower PSNR along with weaker PESQ score.
+Table 5. FewSound results with different encoders.
 Encoder MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
 SNAC 0.007 27.55 1.56 4.03 1.30 0.77 0.32
-Table 6: KAN performance on the 1.5 s samples, with respect to the encoding length. The results are averaged across 10 samples.
+Effect of dataset variability
+We extend the results for our FewSound model by training it on three additional speech datasets - noisy VCTK (Valentini-Botinhao et al., 2016), a version of the VCTK dataset with added noise, LibriSpeech (Panayotov et al., 2015) and LJ Speech (Ito and Johnson, 2017).
+Moreover, it is also evaluated on GTZAN (Tzanetakis and Cook, 2002) and UrbanSound8K (Salamon et al., 2014), which represent music and environmental noise, respectively.
+We follow the same setup as in our main experiment, with one exception—in this experiment, we use a hypernetwork MLP with 41.2M parameters instead of the standard one.
+In case of Librispeech, we train on the train-clean-100 partition and use the test-other partition as the validation set.
+Since PESQ, STOI and CDPAM are suitable only for clean speech datasets, they are omitted from the FewSound model assessment on noisy VCTK, GTZAN and UrbanSound8K.
+All obtained results are presented in Table 6.
+The results indicate that FewSound achieves lower MSE and higher SI-SNR on speech datasets than on music or environmental noise, likely due to encoder limitations in handling non-speech signals.
+Between the KAN and NeRF targets, KAN consistently outperforms in LSD and PESQ across all datasets.
+For SI-SNR and PSNR, performance is comparable, e.g., NeRF excels on LibriSpeech, while KAN leads on LJ Speech.
+Table 6. FewSound performance across different datasets.
+Dataset Target MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+[…]
+VCTK NERF 0.006 28.39 1.50 5.03 1.41 0.79 0.30
+KAN 0.006 28.35 1.46 5.11 1.45 0.79 0.34
+noisy NERF 0.013 24.76 1.42 2.15 - - -
+VCTK KAN 0.013 24.69 1.31 3.31 - - -
+Libri NERF 0.016 23.8 2.42 -2.18 1.26 0.71 0.34
+Speech KAN 0.018 23.47 1.93 -3.21 1.31 0.73 0.31
+LJ NERF 0.013 24.80 1.66 -0.72 1.22 0.74 0.38
+Speech KAN 0.012 25.11 1.51 -0.43 1.24 0.76 0.35
+GTZAN NERF 0.031 21.21 1.49 -4.23 - - -
+KAN 0.028 21.59 1.34 -3.50 - - -
+Urban NERF 0.051 18.94 1.28 -12.93 - - -
+Sound-8k KAN 0.049 19.15 1.22 -11.86 - - -
+4.3. KAN Ablation Study
+Table 7. KAN performance on the 1.5 s samples, with respect to the encoding length. The results are averaged across 10 samples.
 Enc. Len. #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-2 23,016 0.0033 30.7170 2.3927 9.0285 1.4980 0.7556 0.3731
-4 25,704 0.0017 33.7562 1.6014 12.3093 1.9463 0.9040 0.2561
-8 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
-10 33,768 0.0005 41.1152 1.3921 20.0078 3.4947 0.9930 0.1574
-12 36,456 0.0005 42.6504 1.4182 21.6062 3.5176 0.9949 0.1664
-16 41,832 0.0002 46.8576 1.4194 25.9786 3.6256 0.9957 0.1538
-24 52,584 0.0003 45.7863 1.4764 24.9128 3.6033 0.9955 0.1982
-Table 7: KAN results on 1.5 s samples in relation to different number and sizes of hidden layers.
+2 23,016 3.3e-3 30.72 2.39 9.03 1.50 0.76 0.37
+4 25,704 1.7e-3 33.76 1.60 12.31 1.95 0.90 0.26
+8 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
+10 33,768 5e-4 41.12 1.39 20.01 3.49 0.99 0.16
+12 36,456 5e-4 42.65 1.42 21.61 3.52 0.99 0.17
+16 41,832 2e-4 46.86 1.42 25.98 3.63 1.00 0.15
+24 52,584 3e-4 45.79 1.48 24.91 3.60 1.00 0.20
+Table 8. KAN results on 1.5 s samples in relation to different number and sizes of hidden layers.
 Layer Sizes #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-[24, 12, 6] 10,500 0.0014 35.9192 1.4806 14.5340 2.1313 0.9360 0.2977
-[24, 24, 12, 6] 18,564 0.0012 37.6839 1.3589 16.4163 2.7035 0.9778 0.2746
-[48, 12, 6] 19,908 0.0010 38.5736 1.3589 17.3090 2.8569 0.9795 0.2679
-[48, 12, 12, 6] 21,924 0.0010 38.2946 1.3557 17.0775 2.9597 0.9810 0.2618
-[48, 16, 8] 23,408 0.0009 38.7324 1.3347 17.5094 3.0476 0.9845 0.2451
-[48, 24, 12] 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
-[96, 48, 24] 102,480 0.0003 43.6428 1.1715 22.5679 4.0203 0.9964 0.1072
-We further analyze the influence of various parameters’ values on the KAN’s performance.
-In the ablation study, experiments are also conducted on audio samples of 1.5 seconds and 3 seconds, with each subset limited to only 10 recordings.
-Due to the page limit, we include selected results for 1.5 s audio.
-More parameters and analysis of 3 s audio samples can be found in Appendix A.
-During our study we keep the following base parameters’ values, changing only the analyzed components: encoding length = 8, grid size = 10, spline order = 2, layer sizes = [48, 24, 12], scale noise = 0.1, scale base = 1, scale spline = 1, learning rate = 5e-3 and enable scale spline = True, which controls whether we include the learnable spline weigths matrix $w_{s}$.
+[24, 12, 6] 10,500 1.4e-3 35.92 1.48 14.53 2.13 0.94 0.30
+[24, 24, 12, 6] 18,564 1.2e-3 37.68 1.36 16.42 2.70 0.98 0.27
+[48, 12, 6] 19,908 1e-3 38.57 1.36 17.31 2.86 0.98 0.27
+[48, 12, 12, 6] 21,924 1e-3 38.29 1.36 17.08 2.96 0.98 0.26
+[48, 16, 8] 23,408 9e-4 38.73 1.33 17.51 3.05 0.98 0.25
+[48, 24, 12] 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
+[96, 48, 24] 102,480 3e-4 43.64 1.17 22.57 4.02 1.00 0.11
+Next, we analyze the influence of various parameters’ values on the KAN’s performance.
+In this ablation study, experiments are also conducted on audio samples of 1.5 seconds and 3 seconds, with each subset limited to only 10 recordings.
+During our study we keep the following base parameters’ values, changing only the analyzed components: encoding length = 8, grid size = 10, spline order = 2, layer sizes = [48, 24, 12], learning rate = 5e-3 and enable scale spline = True, which controls whether we include the learnable spline weigths matrix $w_{s}$.
 Theoretically, these weigths are redundant, as they could be absorbed into the spline itself.
-Table 6 illustrates the impact of the length of the encoding (size of the positional encoding) that directly influences the size of networks.
+In this section, we focus on the 1.5s samples and the most important parameters.
+The remaining ablation studies can be found in section A.1.
+Table 7 illustrates the impact of the length of the encoding (size of the positional encoding), one of the parameters directly influencing the size of the network.
 Values smaller than the baseline of 8 result in progressively worse performance; meanwhile, increasing the length provides improvement up to value of 16, where the results are the best for all but one metric (LSD).
-Table 7 demonstrates the impact of different hidden layer configurations.
-The results indicate a consistent performance improvement as network capacity increases, with the architecture containing hidden layers of [96, 48, 24] achieving the best performance.
-Decreasing the networks size leads to progressive results degradation.
-Table 8 presents a comparison of the varying spline order values (the degree of the applied spline).
+Table 8 shows that increasing hidden layer capacity consistently improves performance, with the [96, 48, 24] configuration performing best. Smaller architectures yield progressively worse results.
+Furthermore, Table 9 presents a comparison of the varying spline order values (the degree of the applied spline).
 Increasing the spline order allows creation of more complex spline functions; however, this comes at the cost of additional parameters.
@@
 Using smaller or larger spline orders leads to a decline in metric scores.
-Table 9 provides insights into the influence of grid size (number of knots defining the spline) on KAN’s learning process.
+Finally, Table 10 provides insights into the influence of grid size (number of knots defining the spline) on KAN’s learning process.
 Similarly to encoding length and spline order, an increase in grid size leads to a higher number of INR parameters.
@@
 Extra parameters enhance the expressiveness of the INR, improving all metrics except LSD.
-Table 8: KAN spline order (O) for 1.5 s audio recordings
+Table 9. KAN spline order (O) for 1.5 s audio recordings
 O Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
 7 42,180 9e-4 38.91 1.23 17.71 3.21 0.98 0.21
-Table 9: Impact of spline function grid size (G) on 1.5 s reconstructions
+Table 10. Impact of spline function grid size (G) on 1.5 s reconstructions
 G Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
 17 46,620 3e-4 44.78 1.31 23.78 3.81 1.00 0.13
+4.4. Frequency-based analysis
 Spectrograms
@@
 Across all methods, the harmonic structures in the low-frequency range are generally well-preserved, reflecting their ability to capture fundamental features of the audio signal.
-However, significant variations emerge in the representation of high-frequency components, where most INRs lowers the intensity of the components.
+However, significant variations emerge in the representation of high-frequency components, where most INRs lower the intensity of the components.
 It is particularly visible with FINER.
@@
 Finally, one can note the lower average intensity in the reconstructed audio.
-Figure 2: The comparison of spectrograms. We compare the original sample with the waveforms reconstructed using different INR approaches.
-5 Conclusion
+Figure 3. The comparison of spectrograms. We compare the original sample (GT) with the waveforms reconstructed using different INR approaches.
+Spectral Analysis
+To demonstrate KAN’s improvements over previous INRs in high-frequency information of the reconstruction, we designed an additional experiment.
+We can treat the absolute values of the FFT of the audio signal as probability distribution of the base frequencies.
+Then, we can compare distributions of the ground truth and the reconstruction using Wasserstein distance.
+By using $[0,1,4,...(len(audio)-1)^{2}]$ as the support of the sample instead of intuitive $[0,1,2,...len(audio)-1]$, we can make the metric more biased towards high frequencies.
+Using this metric in our standard experimental setting for INRs, we obtain the following results, presented in Table 11.
+They indicate that KAN achieves the lowest score of 79.72 among all baseline methods.
+The second-best outcome is obtained by WIRE, while RFF performs the worst.
+Table 11.
+Wasserstein Distance (WD) of reconstruction and ground truth frequency distributions with the high-frequency biased support.
+For readability purposes the actual values are scaled down by a constant factor of 32768, that is the length of the sample (lower score is better).
+NERF SIREN FINER WIRE RFF KAN
+WD 184.79 218.54 205.75 131.56 362.98 79.72
+5. Conclusion
 This work demonstrates the effectiveness of Kolmogorov-Arnold Networks (KANs) as implicit neural representations for audio signals, offering high reconstruction accuracy when integrated with hypernetwork-based architectures.
 By evaluating KANs against classical INR models like NERF, FINER, WIRE, RFF, and SIREN, we highlight their potential as a foundational approach for efficient audio modeling, resulting in top-2 mean in 5 out of 7 considered quality metrics for 1.5 s audio, and top-2 in all 7 for 3 s audio.
-Additionally, by integrating KAN into hypernetwork frameworks like FewSound, performance improves significantly, surpassing HyperSound with a 33.3% reduction in MSE, a 60.87% increase in SI-SNR, and an 8.66% in PESQ.
+Additionally, we introduce a new hypernetwork framework called FewSound, and show that by integrating KAN performance improves significantly, surpassing HyperSound with a 33.3% reduction in MSE, a 60.87% increase in SI-SNR, and an 8.66% in PESQ.
 The results suggest that KANs can serve as a powerful tool for audio representation, enabling adaptive processing across various datasets and tasks.
-Since our experiments were conducted at a fixed sampling rate of 22,050 Hz, the impact of varying sampling strategies requires investigation in future work.
-Furthermore, integration with modern neural audio codecs and speech encoder systems presented in Tab 5 might be further expanded, in particular when investigating multilingual approach.
-The evaluation was restricted to English datasets due their prevalence.
-Additionally, the datasets lacked complex auditory scenarios, such as multi-speaker dialogues, emotional speech, and diverse acoustic scenes, which might be an interesting extensions to the research.
-It is important to note, that while KANs demonstrated efficiency in certain tasks, their computational demands may grow significantly for longer sequences or large-scale applications.
-Future research should extend KAN-based models to multilingual datasets and diverse linguistic settings to enhance their generalization.
-Investigating the impact of sample length on performance could provide deeper insights into the balance between computational efficiency and fidelity.
-Architectural refinements may reduce resource requirements while maintaining accuracy, making KANs more scalable for practical deployment.
-Exploring their integration with multimodal frameworks, such as facial expressions or gestures, could enable advanced audio-visual synthesis and applications like avatar generation, an interesting path given hypernetworks already being used in visual modality (Przewiezlikowski et al., 2024; Kania et al., 2024).
-Additionally, developing interpretable models would improve transparency and understanding of KANs’ decision-making processes, which would increase trust in real-world applications and meeting legal requirements for practical usage.
-Finally, optimizing compatibility with a broader range of audio encoding systems could significantly enhance reconstruction quality, expanding the practical utility of KAN-based audio representations.
-Appendix A Extended ablation study
-At the beginning of this section we conclude the ablation study for 1.5 s audio samples by analyzing the influence of learning rate, standalone scale spline flag, scale base factor, scale noise factor and scale spline factor on KAN’s performance.
-Table 10: Effect of various learning rate values on 1.5 s recordings
-Learning rate Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
-0.01 31,080 0.0007 40.5596 1.3698 19.4728 3.2799 0.9877 0.1919
-0.007 31,080 0.0007 40.5170 1.3284 19.3607 3.3883 0.9908 0.1927
-0.006 31,080 0.0007 40.0159 1.3051 18.9209 3.3532 0.9911 0.1770
-0.005 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
-0.001 31,080 0.0011 37.8920 1.3626 16.6339 2.5975 0.9711 0.2314
-Table 10 shows the comparison of learning rates.
-Generally, KAN performs well and stably with higher learning rate setting.
-A learning rate of 0.001 yields the worst scores across all entries in the table.
-A learning rate of 0.01 achieves the best results in the PSNR and Si-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores.
-Table 11 shows the comparison of performance of KAN depending on whether the enable scale spline flag is set to False or True.
-The former results in skipping the theoretically redundant weigths.
-However, it leads to a drop in all metric scores except MSE, with perceptual metrics being affected the most.
-Table 11: Enable scale spline flag ablation for 1.5 s audio source
-Enable Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
-False 28,860 0.0007 39.4067 1.5253 18.2321 2.7388 0.9653 0.2771
-True 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
-At the end of this paragraph, we consider hyperparameters related to KAN initialization.
-The weight matrices $w_{b}$ and $w_{s}$ are initialized according to Kaiming uniform schema - the vaules are drawn from $\mathcal{U}(-k,k)$ distrubution, where
-$k=\sqrt{\frac{6}{(1+5a^{2})\cdot fan\_in}}$
-and $a$ is equal to scale base or scale spline for $w_{b}$ and $w_{s}$ respectively.
-The $c_{i}$s parameters are initialized by sampling a noise from $\mathcal{N}(0,\sigma^{2})$ distribution, and solving an instance of the least squares problem, so that $spline(x)$ for a certain $x$ is as close to the sampled noise as possible.
-The scale noise parameter is the standard deviation $\sigma$ of the noise distribution.
-Tables 12 and 13 show the ablation results for scale spline and scale base parameters respectively.
-In case of the first one, 0.5 seems to be the most optimal choice with the highest score for all metrics excluding LSD, where optimal choice is 1, and STOI, where the difference between the best scores is negligible.
-In case of scale spline the choice of the best value varies between metrics.
-Table 14 shows the results for the study of scale noise hyperparameter, with the lowest tested value of 0.025 attaining the best results in most of the metrics.
-Table 12: KAN performance on 1.5 s, with respect to the scale base values.
-Scale Base Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
-0.25 31,080 0.0008 39.9112 1.3386 18.7366 3.3149 0.9888 0.1872
-0.5 31,080 0.0006 41.0193 1.3111 19.8693 3.4322 0.9904 0.1875
-1 31,080 0.0006 40.8162 1.2895 19.6861 3.3821 0.9900 0.1930
-2 31,080 0.0006 40.5262 1.3129 19.4204 3.2141 0.9906 0.2013
-4 31,080 0.0008 39.5773 1.3195 18.4513 3.2608 0.9876 0.2062
-Table 13: KAN performance on 1.5 s, with respect to the scale spline values.
-Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
-0.25 31,080 0.0006 40.6718 1.3196 19.5770 3.5030 0.9917 0.1932
-0.5 31,080 0.0005 40.9010 1.3057 19.8454 3.4086 0.9915 0.1895
-1 31,080 0.0007 40.5384 1.2895 19.4369 3.4281 0.9915 0.1872
-2 31,080 0.0006 41.5030 1.2973 20.3684 3.3479 0.9914 0.1815
-4 31,080 0.0006 40.8259 1.2981 19.7534 3.3932 0.9903 0.1887
-Table 14: KAN performance on 1.5 s, with respect to the scale noise values.
-KAN Scale Noise Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
-0.4 31,080 0.0007 40.9740 1.3010 19.8608 3.4433 0.9910 0.1896
-0.2 31,080 0.0007 40.2776 1.2849 19.1855 3.4280 0.9915 0.1796
-0.1 31,080 0.0007 40.5541 1.2796 19.4436 3.4576 0.9913 0.1897
-0.05 31,080 0.0007 40.0844 1.2864 18.9432 3.4004 0.9909 0.1846
-0.025 31,080 0.0006 41.0045 1.2878 19.9189 3.4605 0.9897 0.1825
-Performance analysis of KAN on 3 s sounds
-Next, we present the ablation study for 3 s recordings, considering the same hyperparameters as used for the 1.5 s samples.
+As the experiments used a fixed 22,050 Hz sampling rate, exploring alternative sampling strategies remains a future work.
+Integration with modern neural codecs and speech encoders (Tab.
+5) could be extended, particularly for multilingual scenarios.
+Moreover, our evaluation focused on English due to dataset availability and lacked complex auditory conditions (e.g., multi-speaker, emotional speech, diverse scenes), suggesting avenues for future research.
+It is important to note that while KANs demonstrated efficiency in certain tasks, their computational demands may grow significantly for longer sequences or large-scale applications.
+Future research should extend KAN-based models to multilingual datasets and varied linguistic contexts to improve generalization.
+Analyzing the effect of sample length may offer insight into balancing fidelity and computational efficiency.
+Architectural optimizations could reduce resource usage while preserving accuracy, enhancing scalability.
+Integrating KANs with multimodal frameworks, e.g., facial expressions or gestures could enable advanced audio-visual synthesis, such as avatar generation, especially given the use of hypernetworks in visual domains (Przewięźlikowski et al., 2024; Kania et al., 2024).
+Developing interpretable models would improve transparency and trust, supporting real-world deployment and legal compliance.
+Lastly, enhancing compatibility with diverse audio encoding systems could boost reconstruction quality and expand KANs’ practical utility.
+Appendix A Additional experiments
+A.1. Extended ablation study
+At the beginning of this section we conclude the ablation study (presented initially in Sect.
+4.3).
+We show how KAN performance on 3 s audio samples is influenced by encoding length, layer sizes, spline order and  spline function grid size parameters.
+We consider the same hyperparameters as used for the 1.5 s experiments.
 Table 12 shows the impact of different encoding lengths on 3 s utterances.
@@
 The results are much less conclusive than in case of 1.5 s, where length 16 was clearly optimal across most metrics.
-However, the majority of metrics (MSE, PSNR, SI-SNR and PESQ, STOI) are the best for encoding lengths of 24 and 16 respectively.
-Table 15: KAN performance on the 3 s samples, with respect to the encoding length.
-Encoding Length Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+However, the majority of metrics (MSE, PSNR, SI-SNR and PESQ, STOI) are the best for encoding lengths of 24 and 16, respectively.
+Table 12. KAN performance on the 3 s samples, with respect to the encoding length.
+Encoding #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-2 23,016 0.0036 30.4569 3.3098 4.2598 1.2121 0.6627 0.4217
-4 25,704 0.0011 35.4080 2.0309 10.3129 1.5506 0.8204 0.2926
-8 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
-10 33,768 0.0006 39.3990 1.3152 14.5468 2.8272 0.9779 0.3795
-12 36,456 0.0005 39.6602 1.3722 14.7926 2.7818 0.9799 0.3730
-16 41,832 0.0004 40.7796 1.4383 15.9502 2.9095 0.9820 0.3516
-24 52,584 0.0003 41.9264 1.6033 17.1797 2.8385 0.9816 0.3429
+2 23,016 3.6e-3 30.46 3.31 4.26 1.21 0.66 0.42
+4 25,704 1.1e-3 35.41 2.03 10.31 1.55 0.82 0.29
+8 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
+10 33,768 6e-4 39.40 1.32 14.55 2.83 0.97 0.38
+12 36,456 5e-4 39.66 1.37 14.79 2.78 0.98 0.37
+16 41,832 4e-4 40.78 1.44 15.95 2.91 0.98 0.35
+24 52,584 3e-4 41.93 1.60 17.18 2.84 0.98 0.34
 Table 13 shows how sizes and depths of hidden layers influence results on 3 s samples.
 Similarly to 1.5 s (Table 8) the best performance is achieved by the largest network composed of [96, 48, 24] layers.
-The performance, however does not improve with the increasing size as gradually as for 1.5 s samples.
-Table 16: KAN results on 3 s samples in relation to different number and sizes of hidden layers.
-Layer Sizes Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+The performance, however, does not improve with the increasing size as gradually as for 1.5 s samples.
+Table 13. KAN results on 3 s samples in relation to different number and sizes of hidden layers.
+Layer Sizes #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-[24, 12, 6] 10,500 0.0007 37.5744 1.4904 12.6576 1.7192 0.8912 0.3229
-[24, 24, 12, 6] 18,564 0.0006 38.4196 1.3828 13.5231 1.9078 0.9393 0.3847
-[48, 12, 6] 19,908 0.0008 38.0106 1.3325 13.0259 2.1609 0.9384 0.3734
-[48, 12, 12, 6] 21,924 0.0006 38.3097 1.3430 13.3983 2.1817 0.9517 0.3886
-[48, 16, 8] 23,408 0.0007 38.2895 1.3119 13.3530 2.2985 0.9562 0.3756
-[48, 24, 12] 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
-[96, 48, 24] 102,480 0.0005 40.9599 1.1792 16.1224 3.5224 0.9898 0.1594
+[24, 12, 6] 10,500 7e-4 37.57 1.49 12.66 1.72 0.89 0.32
+[24, 24, 12, 6] 18,564 6e-4 38.42 1.38 13.52 1.91 0.94 0.38
+[48, 12, 6] 19,908 8e-4 38.01 1.33 13.03 2.16 0.94 0.37
+[48, 12, 12, 6] 21,924 6e-4 38.31 1.34 13.40 2.18 0.95 0.39
+[48, 16, 8] 23,408 7e-4 38.29 1.31 13.35 2.30 0.96 0.38
+[48, 24, 12] 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
+[96, 48, 24] 102,480 5e-4 40.96 1.18 16.12 3.52 0.99 0.16
 In comparison to the 1.5 s case, where a spline order of size 3 is clearly the optimal hyperparameter setting, its advantage is less evident in Table 14.
-KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI columns.
-Table 17: KAN effectiveness in terms of diverse spline order sizes and 3 s long input
-Spline Order Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI metrics.
+Table 14. KAN effectiveness in terms of diverse spline order sizes and 3 s long input
+Spline #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-1 28,860 0.0006 38.5038 1.4711 13.6106 2.2903 0.9540 0.3299
-2 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
-3 33,300 0.0008 38.6639 1.2540 13.6820 2.5639 0.9704 0.3578
-5 37,740 0.0006 39.0929 1.2663 14.1943 2.4499 0.9661 0.3376
-7 42,180 0.0006 38.8824 1.3175 13.9657 2.1987 0.9573 0.3378
-In table 18, table 19 and table 20, which correspond respectively to the ablations of grid size, learning rate, and enable scale spline flag on 3 s speeches, we observe similar conclusions and tendencies as in the 1.5s scenario.
-Described in Table 9 (Sect.
-4.2), and Tables 10, 11(Appendix A), respectively.
-Table 18: KAN effectiveness in terms of diverse grid size and 3 s long input
-Grid size Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+1 28,860 6e-4 38.50 1.47 13.61 2.29 0.95 0.33
+2 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
+3 33,300 8e-4 38.66 1.25 13.68 2.56 0.97 0.36
+5 37,740 6e-4 39.09 1.27 14.19 2.45 0.97 0.34
+7 42,180 6e-4 38.88 1.32 13.97 2.20 0.96 0.34
+The results of the grid size parameter (displayed in Table 15) provide similar conclusions and tendencies as in the 1.5 s scenario (Table 10) — increasing the grid size leads to the improvement in all metrics but LSD.
+Table 15. KAN effectiveness in terms of diverse grid size and 3 s long input
+Grid size #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-1 11,100 0.0029 31.3359 2.0005 5.4555 1.3828 0.7652 0.3998
-5 19,980 0.0007 38.1135 1.4668 13.2103 1.9651 0.9160 0.2399
-10 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
-12 35,520 0.0005 39.4903 1.2745 14.6419 2.6706 0.9739 0.3204
-17 46,620 0.0006 39.7273 1.3224 14.8089 3.0683 0.9734 0.2646
-Table 19: Effect of various learning rate values on 3 s recordings
-Learning rate Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+1 11,100 2.9e-3 31.34 2.00 5.46 1.38 0.77 0.40
+5 19,980 7e-4 38.11 1.47 13.21 1.97 0.92 0.24
+10 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
+12 35,520 5e-4 39.49 1.27 14.64 2.67 0.97 0.32
+17 46,620 6e-4 39.73 1.32 14.81 3.07 0.97 0.27
+Table 16. Effect of various learning rate values on 1.5 s and 3 s recordings.
+Learning #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-0.01 31,080 0.0006 39.4961 1.3238 14.6358 2.5597 0.9692 0.3940
-0.007 31,080 0.0006 39.2295 1.2804 14.3368 2.5669 0.9707 0.3685
-0.006 31,080 0.0006 39.1037 1.2781 14.2218 2.5516 0.9700 0.3633
-0.005 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
-0.001 31,080 0.0007 37.7570 1.4674 12.8016 1.9302 0.9261 0.3020
-Table 20: Enable scale spline flag ablation for 3 s audio source
-Enable Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+Sample Length: 1.5 s
+0.01 31,080 7e-4 40.56 1.37 19.47 3.28 0.99 0.19
+0.007 31,080 7e-4 40.52 1.33 19.36 3.39 0.99 0.19
+0.006 31,080 7e-4 40.02 1.31 18.92 3.35 0.99 0.18
+0.005 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
+0.001 31,080 1.1e-3 37.89 1.36 16.63 2.60 0.97 0.23
+Sample Length: 3 s
+0.01 31,080 6e-4 39.50 1.32 14.64 2.56 0.97 0.39
+0.007 31,080 6e-4 39.23 1.28 14.34 2.57 0.97 0.37
+0.006 31,080 6e-4 39.10 1.28 14.22 2.55 0.97 0.36
+0.005 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
+0.001 31,080 7e-4 37.76 1.47 12.80 1.93 0.93 0.30
+At the end of this paragraph, we include ablation studies for the remaining hyperparameters.
+Table 16 shows the comparison of learning rates.
+A value of 0.001 yields the worst scores across all entries in the table, with the exception of the CDPAM metric.
+A learning rate of 0.01 achieves the best results in the PSNR and SI-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores.
+Generally, KAN performs well and stably with higher learning rate setting.
+It can be considered its advantage over networks like SIREN, which are sensitive to the choice of the learning rate and their learning process can collapse when too high values are used.
+Table 17 shows the comparison of performance of KAN depending on whether the enable scale spline flag is set to False or True.
+The former results in skipping the theoretically redundant weights.
+However, it leads to a drop in all metric scores except MSE, with perceptual metrics being affected the most.
+Table 17. Effect of enabling the scale spline flag for 1.5 s and 3 s recordings.
+Scale #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
 […]
-False 28,860 0.0007 38.0986 1.4767 13.2310 2.1969 0.9360 0.4218
-True 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
-Finally, in Tables 21, 22, and 23 one can notice that the scale base and the scale spline hyperparameters do not significantly influence the performance.
-While manipulating scale base, all metrics except for CDPAM vary by less than 9% (the mentioned CDPAM under 16%).
-Moreover, apart from MSE that almost does not change, the changes in metric scores are not monotonic with respect to the changes in parameter values.
-Table 21: Evaluation of different scale base factors on 3 s audio samples
-Scale Base Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+Sample Length: 1.5 s
+False 28,860 7e-4 39.41 1.53 18.23 2.74 0.97 0.28
+True 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
+Sample Length: 3 s
+False 28,860 7e-4 38.10 1.48 13.23 2.20 0.94 0.42
+True 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
+A.2. INR
+To show that KAN performs well across diverse types of audio, we train INRs on samples from the two aforementioned datasets - GTZAN (Tzanetakis and Cook, 2002) and UrbanSound8K (Salamon et al., 2014).
+We chose 100 samples for each dataset, 10 samples per genre in case of GTZAN and 10 samples per split in case of UrbanSound8K.
+The recordings were pre-processed in an analogous way to our main experiment - resampled to 22.05 kHz and cropped to 32,768 samples, resulting in audio of length 1.5 s.
+The results are presented in Tables 18 and 19.
+While FINER achieves the best PSNR and SI-SNR, it is the worst with respect to LSD.
+WIRE shows the best overall performance when all metrics are taken into account, followed by SIREN and KAN, both achieving similar results.
+Table 18. Reconstruction metrics for GTZAN dataset. We report mean and standard deviation across 100 samples, 10 samples per genre.
+INR (# params) MSE PSNR LSD SI-SNR
 […]
-0.25 31,080 0.0006 39.0743 1.2804 14.1621 2.4855 0.9654 0.3574
-0.5 31,080 0.0006 39.1813 1.2744 14.2891 2.5813 0.9714 0.3485
-1 31,080 0.0006 38.9647 1.2643 14.0714 2.5415 0.9693 0.3573
-2 31,080 0.0006 38.8621 1.2980 13.9575 2.5120 0.9654 0.3932
-4 31,080 0.0005 39.7455 1.3119 14.9256 2.3783 0.9612 0.4040
-Table 22: Evaluation of different scale spline factors on 3 s audio samples
-Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
+NERF (33,261) 3.6e-3 $\pm$ 4.5e-3 32.70 $\pm$ 5.12 1.21 $\pm$ 0.42 13.16 $\pm$ 5.05
+SIREN (31,191) 3.7e-3 $\pm$ 2.5e-2 37.13 $\pm$ 5.18 1.10 $\pm$ 0.38 17.05 $\pm$ 8.37
+FINER (31,191) 1.0e-3 $\pm$ 2.5e-3 40.89 $\pm$ 5.60 1.23 $\pm$ 0.53 21.36 $\pm$ 4.86
+WIRE (31,191) 2.0e-3 $\pm$ 6.2e-3 37.13 $\pm$ 5.87 1.03 $\pm$ 0.38 17.50 $\pm$ 5.59
+RFF (39,201) 4.9e-3 $\pm$ 2.1e-2 32.94 $\pm$ 4.86 1.25 $\pm$ 0.42 13.63 $\pm$ 4.89
+KAN (33,768) 1.6e-3 $\pm$ 2.8e-3 36.98 $\pm$ 5.78 1.14 $\pm$ 0.45 17.46 $\pm$ 5.87
+Table 19. Reconstruction metrics for UrbanSound8K. We report mean and standard deviation across 100 samples, 10 samples per split.
+INR (# params) MSE PSNR LSD SI-SNR
 […]
-0.25 31,080 0.0006 39.3700 1.2761 14.4961 2.6388 0.9713 0.3758
-0.5 31,080 0.0006 38.9780 1.2698 14.0723 2.5493 0.9689 0.3678
-1 31,080 0.0006 39.0889 1.2654 14.2169 2.4618 0.9643 0.3667
-2 31,080 0.0006 38.9006 1.2735 13.9963 2.5288 0.9696 0.3663
-4 31,080 0.0006 38.9421 1.2740 14.0168 2.4376 0.9666 0.3448
-Table 23: Evaluation of different scale noise factors on 3 s audio samples
-Scale Noise Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
-0.4 31,080 0.0006 39.0755 1.2712 14.1628 2.5252 0.9659 0.3711
-0.2 31,080 0.0006 39.1540 1.2631 14.2794 2.5684 0.9687 0.3604
-0.1 31,080 0.0006 39.0749 1.2624 14.1858 2.5053 0.9698 0.3488
-0.05 31,080 0.0006 39.0345 1.2699 14.1197 2.5075 0.9680 0.3601
-0.025 31,080 0.0006 39.1100 1.2706 14.2064 2.5701 0.9679 0.3613
+NERF (33,261) 3e-3 $\pm$ 4.9e-3 34.23 $\pm$ 5.35 1.25 $\pm$ 0.52 14.81 $\pm$ 6.33
+SIREN (31,191) 1.8e-3 $\pm$ 4.8e-3 37.77 $\pm$ 5.79 1.14 $\pm$ 0.46 17.15 $\pm$ 12.58
+FINER (31,191) 7.1e-4 $\pm$ 1.3e-3 40.46 $\pm$ 5.07 1.27 $\pm$ 0.52 20.98 $\pm$ 4.23
+WIRE (31,191) 1.6e-3 $\pm$ 2.7e-3 38.28 $\pm$ 6.74 1.11 $\pm$ 0.45 18.65 $\pm$ 7.49
+RFF (39,201) 4.1e-3 $\pm$ 7.2e-3 33.47 $\pm$ 6.05 1.23 $\pm$ 0.51 14.09 $\pm$ 7.41
+KAN (33,768) 1.6e-3 $\pm$ 2.1e-3 36.92 $\pm$ 6.02 1.15 $\pm$ 0.49 17.37 $\pm$ 6.15
+GenAI Usage Disclosure
+The authors declare that generative AI tools were not used in the preparation of the manuscript or the development of the corresponding source code, apart from software tools for the evaluation of spelling and grammatical correctness.

### v3

@@
-A Hypernetwork-Based Approach to KAN Representation of Audio Signals
+As Good as It KAN Get: High-Fidelity Audio Representation
 Abstract.
@@
 representation learning, audio representation, INR, KAN
-conference: ; isbn: ;ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Learning paradigmsccs: Applied computing Sound and music computing
+journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761405isbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Learning paradigmsccs: Applied computing Sound and music computing
 Figure 1.
@@
 We propose to analyze sounds using KAN (Kolmogorov-Arnold Network), which uses trainable activation functions instead of classical layers.
-Such an operation can be seen as approximating signals by a linear combination of basis functions, bearing a similarity to the FFT or wavelets.
+Such an operation can be seen as approximating signals by a linear combination of basis functions, bearing a similarity to the FFT or wavelets.:
 1. Introduction
@@
 In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data.
-HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes raw input sound and produces weight for INR-based sound representation.
+HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes raw input sound, and produces weight for INR-based sound representation.
 INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes.
-Siamese SIREN (Lanzendörfer and Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
-The recently proposed KAN (Kolmogorov-Arnold Network) (Liu et al., 2024a) replaces traditional fully connected layers with activation functions.
+Siamese SIREN (Lanzendörfer and Wattenhofer, 2023) is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
+The recently proposed KAN (Kolmogorov-Arnold Network) (Liu et al., 2024a) replaces traditional fully connected layers with learnable activation functions based on splines.
+Any function f (e.g., an audio representation of the magnitude over time) can be approximated by splines with $\sup\|\text{f-spline}\|=\mathrm{O}(h^{4})$ (Hall and Meyer, 1976).
+KANs inherit this theoretical precision as a linear combination of splines, resulting in a polynomial bound.
+This indicates potential effectiveness for audio reconstruction.
 While KAN shows potential, it faces limitations with high-dimensional data.
 However, its effectiveness in low-dimensional settings and successful application to complex function regression tasks make it a promising candidate for audio INR.
+In our work we examine six INR approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a).
 This paper shows that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm.
 For single audio INR, KAN trains more rapidly and achieves better results.
-Setting up the conditions for a KAN is not immediately straightforward.
-We show that hypernetworks can generate KANs.
-In practice, we can directly produce KAN parameters or use updates for the universal KAN-based architecture.
-Hence, we propose two methods, a classical HyperSound (Szatkowski et al., 2023) with a KAN target network.
-The second uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR, which is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017).
-As there is no few-shot model for audio signals, we introduce FewSound  [^1] The source code for the proposed method can be found at the following link: https://github.com/gmum/fewsound.
-It takes raw sound on input and uses a hypernetwork to generate updates to the universal weights of INR dedicated to encoding sound.
-In our work, we examine six approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a).
-In the model, we simultaneously train the hypernetwork weights and the universal weights.
+We also show that KANs can be generated by hypernetworks.
+In practice, we can directly produce KAN parameters or use updates for a universal KAN-based architecture.
+Hence, we propose two methods.
+The first one is classical HyperSound (Szatkowski et al., 2023) with a KAN target network, while the second one uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR.
+As there is no few-shot model for audio signals, we introduce FewSound.
+[^1] The source code for the proposed method can be found at the following link: https://github.com/gmum/fewsound It takes raw sound on input and uses a hypernetwork to generate updates to the universal weights of an INR dedicated to encoding sound.
+FewSound is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017).
 Thus, two settings show that KAN can be effectively modeled by hypernetworks.
-Our study demonstrates that KAN can be viewed as INR of sound.
-Such models yield favorable outcomes when dealing with individual audio signals and excel in large Hypernetwork-based models trained on large datasets.
-We use KANs due to their spline-based activations.
-Since any function f (e.g., an audio representation of the magnitude over time) can be approximated by splines with $\sup\|f\_spline\|=\mathrm{O}(h^{4})$ (Hall and Meyer, 1976), KANs inherit this theoretical precision as a linear combination of splines, resulting in a polynomial bound.
-This suggests that KANs could be effective for audio reconstruction.
 Key contributions of our research include:
   • introducing KAN-based sound representation and showing its efficacy,
-  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN network and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
+  • introducing FewSound, a hypernetwork-based meta-learning method, which is used with KAN networks and improves over previous hypernetwork approach HyperSound (Szatkowski et al., 2023),
   • comparing KAN represenation with other INRs,
@@
 However, due to the ReLU’s second derivative being zero, they cannot capture the fine details of the information located in higher-order derivatives leading to worse performance.
-To counteract this, methods such as SIREN (Sitzmann et al., 2020) have been developed, which use a periodic activation function the sine in place of ReLU.
+To counteract this, methods as SIREN (Sitzmann et al., 2020) have been developed, which use a periodic activation function in place of ReLU.
 The i-th layer of the network is defined as: $\phi_{i}(x_{i})=\sin(\omega_{i}W_{i}^{T}x_{i}+b_{i}),$ where $x_{i}$ is layer’s input, while $W_{i}$ and $b_{i}$ are, respectively, weights and biases.
@@
 Inspired by harmonic analysis, WIRE uses complex Gabor wavelet $\psi$ as its nonlinearity activation: $\psi(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$.
-INRs can suffer from a phenomenon called spectral bias (Rahaman et al., 2019) – the preference of MLPs to learn low-frequency functions and ignore high-frequency noise, which can lead to performance degradation when high-frequency components of the input carry meaningful information (Gorji et al., 2023).
-A common way to overcome spectral bias is by introducing high-frequency embeddings (Mildenhall et al., 2020; Tancik et al., 2020) that modify the space over which the MLP operates, altering the frequency of its output function.
-FINER (Liu et al., 2024b) is an MLP-based architecture designed to flexibly tune the spectral bias.
-It uses a variable-periodic activation $a(x)=sin(\omega_{0}\lvert x+1\rvert x)$ and a specialized bias initialization schema in its first layer.
-NeRF (Neural Radiance Field) (Mildenhall et al., 2020) models scenes via a 5D input—spatial location $(x,\,y,\,z)$ and viewing direction $(\theta,\,\phi)$—with an MLP outputting color $c=(r,\,g,\,b)$ and density $\sigma$, enabling 3D reconstruction from sparse views.
-Figure 2.
-The overview of FewSound architecture.
-The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights.
-The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$.
-The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
+INRs are prone to spectral bias (Rahaman et al., 2019), where MLPs favor low-frequency functions and neglect high-frequency signals, reducing performance when such components are informative (Gorji et al., 2023).
+This is commonly mitigated with high-frequency embeddings (Mildenhall et al., 2020; Tancik et al., 2020), which reshape the input space and adjust output frequencies.
+FINER (Liu et al., 2024b), an MLP-based architecture, flexibly tunes spectral bias via a variable-periodic activation $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$ and specialized first-layer bias initialization.
+NeRF (Neural Radiance Field) (Mildenhall et al., 2020) represents scenes with a 5D input: spatial location (x,y,z)(x,\,y,\,z) and viewing direction $(\theta,\,\phi)$, and an MLP that outputs color c=(r,g,b)c=(r,\,g,\,b) and density $\sigma$, enabling 3D reconstruction from sparse views.
 Sound representation
-The audio encoders we rely on are classified into two main types: neural speech codecs and those employed in automatic speech recognition (ASR) or ASR-adjacent solutions.
+In our research, we use learned representations from deep neural networks, which outperform traditional hand-crafted features (Baevski et al., 2020).
+These representations fall into 2 categories: acoustic embeddings from neural speech codecs (Défossez et al., 2023), and semantic embeddings (Hsu et al., 2021).
+We focus on representations derived from models designed for ASR or related tasks.
 SoundStream (Zeghidour et al., 2021) is a popular neural audio codec based on the encoder-decoder architecture.
@@
 The encoder is based on the linear-spectrogram input that extracts a feature vector composed of fundamental frequency, aperiodicity and loudness of the signal.
+Figure 2.
+The overview of FewSound architecture.
+The input raw sound is processed by encoding network $E(\cdot)$ and, together with encoded universal weights, passed to the fully connected Hypernetwork $H(\cdot)$.
+The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$.
+Finally, the target network $f_{\theta^{\prime}}$ maps time samples to sound amplitudes specific to the considered task.
+:
 KAN (Kolmogorov-Arnold Network)
-KANs (Liu et al., 2024a) is a type of neural network architecture used as an alternative to MLPs, instead of using fixed activation functions.
+KAN (Liu et al., 2024a) is a type of neural network architecture used as an alternative to MLPs, instead of using fixed activation functions.
 KANs utilize learnable activation functions placed on edges rather than nodes.
-KAN’s learnable functions are parametrized as a spline and serve as ”weights”, whereas the nodes only sum the incoming signals.
+KANs’ learnable functions are parametrized as a spline and serve as ”weights”, whereas the nodes only sum the incoming signals.
 This is described by the following formula:
@@
 where $\Phi_{q}$ is a KAN layer, $\phi_{q,\,p}$ is a single activation function and $x_{p}$ is input feature.
-The authors also introduced an architecture of LANs (learnable activation networks) – the only difference between them and MLP lies in the learning aspect of the activation functions.
-The authors show that a network can be used in implicit neural representation tasks.
+The authors introduced LANs (learnable activation networks), which differ from MLPs only by learning activation functions.
+They demonstrate that such networks can be applied to implicit neural representation tasks.
 Audio INR
@@
 3. KAN in hypernetwork-based architectures
-In this section, we show how KAN works with hypernetworks.
-We use two main approaches.
-First, we use the autoencoder-based solution HyperSound (Szatkowski et al., 2023), where we add KAN as the target model.
-We also examine KAN in a few-shot setting introduced by (Wang et al., 2020).
-To examine such a model, we propose FewSound, which is a realization of (Wang et al., 2020) in the case of KAN and hypernetworks.
+In this section, we demonstrate how KAN integrates with hypernetworks using two main approaches.
+First, we extend the autoencoder-based HyperSound (Szatkowski et al., 2023) by employing KAN as the target model.
+Second, we explore KAN in a few-shot setting introduced by (Wang et al., 2020).
+To this end, we propose FewSound, which adapts the framework of (Wang et al., 2020) to the context of KANs and hypernetworks.
 Hypernetwork
-In the foundational study (Ha et al., 2016) hyper-networks are described as neural models that produce weights for another network tasked with solving a specific problem.
-The researchers aim to decrease the number of trainable parameters by creating a hyper-network with fewer parameters than the target network.
-Drawing a parallel between hypernetworks and generative models, the authors of (Sheikh et al., 2017) employ this technique to generate a variety of target networks that approximate the same function.
+In the seminal work (Ha et al., 2016), Hypernetworks are introduced as neural models that produce the weights of another network dedicated to solving a specific task.
+The approach aims to reduce the number of trainable parameters by employing a hypernetwork with fewer parameters than the target network.
+Building on this concept, numerous studies have emerged.
+Among them, (Sheikh et al., 2017) draws an analogy to generative models, demonstrating how this technique can produce a variety of target networks approximating the same function.
+In the audio INR domain, HyperSound (Szatkowski et al., 2023) leverages hypernetworks to transform raw audio into weights of an INR-based representation.
 Few-shot learning for INR
-A significant limitation of deep learning models is their reliance on extensive datasets.
-In contrast, humans can learn effectively from just a few examples.
-A few-shot learning (Wang et al., 2020) approach was proposed to bridge the gap between deep neural network training and human learning capabilities.
-This strategy treats the data set as consisting of numerous small tasks.
-Few-shot architectures can be quickly adapted to solve new problems in a single step.
-The most popular approach to few-shot learning is MAML (Finn et al., 2017).
-MAML uses a set of universal weights, fine-tuned for specific small tasks through a limited number of gradient updates.
-Due to the model-agnostic nature of the approach, it can be adapted to various models and corresponding learning algorithms.
-The terminology of few-shot learning is often inconsistent due to varying definitions in existing research.
-For a standardized framework, see  (Chen et al., 2019; Wang et al., 2020).
-However, such a nomenclature is dedicated to a classification, regression, or generative model task.
-In the case of the INR-based model, we refer to (Tancik et al., 2021).
-We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
+A key limitation of deep learning models is their dependence on large datasets, whereas humans can learn effectively from just a few examples.
+Few-shot learning (Wang et al., 2020) seeks to bridge this gap by structuring data into numerous small tasks and using architectures capable of rapid adaptation to new problems in a single step.
+The most popular method, MAML (Finn et al., 2017), leverages a set of universal weights that are fine-tuned for specific small tasks through a limited number of gradient updates.
+Its model-agnostic nature allows it to be applied to a wide range of models and learning strategies.
+Nevertheless, further generalization is possible.
+HyperMAML (Przewięźlikowski et al., 2024) achieves this by replacing the gradient-based adaptation step with a Hypernetwork that predicts task-specific weight updates directly from the data.
+The terminology of few-shot learning is often inconsistent due to differing definitions in prior work.
+Standardized frameworks exist (Chen et al., 2019; Wang et al., 2020), but their nomenclature is focused on a classification, regression, or generative task.
+For INR-based models, we follow (Tancik et al., 2021).
+In this setting, we have a dataset $X$ of sound samples and aim to produce an INR $f_{\theta}$ that encodes the audio signals in its weights.
 FewSound
-Now we are ready to present FewSound – a model that uses Hypernetworks to update the INR weights dedicated to sound.
-The core concept of this update mechanism is to use information extracted from raw sound data and universal weights to determine the optimal updates for a particular task.
-This method adjusts the INR’s parameters for entirely different sounds.
-The schema of FewSound is provided in Fig. 2. We aim to reconstruct the training example using the target network. In practice, we run our INR on all steps to reconstruct the input.
-Following our previous analysis of INR in the sound reconstruction task, we consider the parameterized function $f_{\theta}$ to be one of the discussed INRs.
-In addition, we distinguish the trainable encoding network $E(\cdot)$ in our architecture, which transforms raw audio data into low-dimensional representation.
-We investigated various encoder architectures and report their performance in Tab.
+At this stage, we are ready to present FewSound, a model that employs Hypernetworks to adapt universal INR weights specialized for sound.
+As illustrated in Fig.
+2, its key idea lies in integrating the information extracted from raw sound data and universal weights to determine the optimal adjustment for a particular task.
+As a result, our method is able to generate INR’s parameters for entirely different sounds.
+Within the FewSound framework, we distinguish a trainable encoding network $E(\cdot)$ that maps raw audio data to a low-dimensional representation $\mathbf{E}_{\mathcal{S}}$.
+We explore several encoder architectures, including SoundStream, Whisper, SNAC, and Spectrogram Encoder, and report their performance in Tab.
 5.
-The encoded sound is concatenated with universal weights propagated by the universal weight encoder $\mathbf{G}(\cdot)$ and then to the fully connected hypernetwork $H(\cdot)$.
-The hypernetwork then modifies the universal weights: $\theta^{\prime}=\theta+\Delta_{\theta}.$ In contrast to the traditional MAML approach, which uses a gradient-based adaptation step, we propose employing a Hypernetwork to anticipate the values directly from the data.
-Target model parameters are computed via Eq. 1.
+In addition, we define a universal weight encoder $\mathbf{G}(\cdot)$ that compresses universal weights $\theta$ to a representation $\mathbf{E}_{\mathcal{\theta}}$ of the same dimensionality as $\mathbf{E}_{\mathcal{S}}$.
+Both representations are concatenated and subsequently processed by a fully connected hypernetwork $H(\cdot)$, which outputs the adjustment $\Delta_{\theta}$ to the universal weights for the given input task.
+Consequently, the final target model parameters $\theta^{\prime}$ can be expressed as:
 (1) $\theta^{\prime}=\theta+\Delta_{\theta}=\theta+H(\mathbf{E}_{\mathcal{S}},\,\mathbf{E}_{\theta}).$
-While FewSound accepts a fixed-size input (in our case 32768), it can be easily used on audio samples of arbitrary length by reconstructing them piece by piece and smoothing during merging.
+Using these adapted weights, we can reconstruct the training example by running one of the INRs described in Section 2 over all time indices of the input signal.
+Although FewSound accepts a fixed-size input (in our case 32768), it can also be used on audio samples of arbitrary length by reconstructing them piece by piece in an overlapping window scheme and applying smoothing during concatenation.
 Training
-For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$.
-The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$.
-During training, we first sample the batch of sounds $\mathcal{B}$ from the dataset $X$.
-Next, for each task, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1).
-Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
-(2) $\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$
-where $\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\lambda_{t}L_{t}(x,\hat{x})+\lambda_{f}L_{f}(x,\hat{x})$, with $L_{t}$ is L1 loss in temporal space, $L_{f}$ is multi-resolution mel-scale STFT loss in frequency space (Yamamoto et al., 2020) and $\lambda$s are coefficients controlling influence of particular losses.
-The Enconder parameters, universal weight encoder, Hypernetwork, and global parameters $\theta$ represent the meta parameters of the system and are updated with stochastic gradient descent (SGD) by optimizing $\mathcal{L}_{HyperMAML}(f_{\theta})$.
+In the training procedure, we assume that the encoder $E(\cdot)$ is parameterized by $\gamma$ ($E:=E_{\gamma}$), the universal weight encoder $G(\cdot)$ by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$ ($H:=H_{\eta}$).
+During each training iteration, we first sample the batch of sounds $\mathcal{B}$ from the dataset $X$.
+Next, the hypernetwork computes the task-specific update $\Delta_{\theta_{i}}$, yielding updated weights $\theta^{\prime}_{i}$ as given in (1).
+These parameters define the corresponding INR $f_{\theta^{\prime}_{i}}$.
+Finally, we obtain the reconstruction $\hat{x}$ of the input sound $x$ using time steps from [1,,1][-1,,1], and train the system parameters with the loss function:
+(2) $\mathcal{L}_{FewSound{}}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}(f_{\theta^{\prime}_{i}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}(f_{\theta+\Delta_{\theta_{i}}})~,$
+[…]
 Table 1. Reconstruction metrics for 1.5 s and 3 s audio. We report mean and standard deviation across 110 samples, each of a different speaker of VCTK dataset (Veaux et al., 2013).
@@
 INR (# params) MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 NERF (33,261) 8.1e-4 $\pm$ 1.3e-3 39.68 $\pm$ 5.67 1.36 $\pm$ 0.18 17.83 $\pm$ 6.10 3.21 $\pm$ 0.48 0.98 $\pm$ 0.04 0.25 $\pm$ 0.11
@@
 INR (# params) MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 NERF (33,261) 4.6e-4 $\pm$ 5.0e-4 40.83 $\pm$ 4.50 1.47 $\pm$ 0.16 17.54 $\pm$ 4.29 2.54 $\pm$ 0.45 0.94 $\pm$ 0.05 0.30 $\pm$ 0.14
@@
 Both subsets consist of a total of 110 recordings.
-To make a fair comparison, all of the MLP based INRs have the same number of layers and with the same sizes, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models.
-All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov and Hutter, 2019) and the same loss function as in FewSound.
+To make a fair comparison, all of the MLP based INRs have the same number of layers with the same sizes, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models.
+All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov and Hutter, 2019) and the same loss function $\mathcal{L}$ as in Eq.
+2.
 Table 1 shows the results of the experiments.
@@
 All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch.
-We use AdamW optimizer (Loshchilov and Hutter, 2019) and OneCycleLR scheduler (Smith and Topin, 2018).
+We use AdamW optimizer (Loshchilov and Hutter, 2019) and OneCycleLR scheduler (Smith and Topin, 2018b).
 The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is $1\mathrm{e}{-6}$.
@@
 Target MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+Network $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 NERF 0.006 28.07 1.50 4.81 1.38 0.78 0.31
@@
 Encoder MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 SoundStream 0.006 28.07 1.50 4.81 1.38 0.78 0.31
@@
 Dataset Target MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+Net. $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 VCTK NERF 0.006 28.39 1.50 5.03 1.41 0.79 0.30
@@
 Enc. Len. #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 2 23,016 3.3e-3 30.72 2.39 9.03 1.50 0.76 0.37
@@
 Layer Sizes #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 [24, 12, 6] 10,500 1.4e-3 35.92 1.48 14.53 2.13 0.94 0.30
@@
 O Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 1 28,860 9e-4 38.83 1.53 17.59 3.01 0.98 0.27
@@
 G Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 1 11,100 3e-3 30.65 2.00 8.99 1.65 0.80 0.43
@@
 4.4. Frequency-based analysis
+Spectral Analysis
+To demonstrate KAN’s improvements over previous INRs in high-frequency information of the reconstruction, we designed an additional experiment.
+We can treat the absolute values of the FFT of the audio signal as probability distribution of the base frequencies.
+Then, we can compare distributions of the ground truth and the reconstruction using Wasserstein distance.
+By using $0,1,4,...(len(audio)-1)^{2}]$ as the support of the sample instead of intuitive [0,1,2,len(audio)1][0,1,2,...len(audio)-1], we can make the metric more biased towards high frequencies.
+Using this metric in our standard experimental setting for INRs, we obtain the following results, presented in Table 11.
+They indicate that KAN achieves the lowest score of 79.72 among all baseline methods.
+The second-best outcome is obtained by WIRE, while RFF performs the worst.
+Table 11.
+Wasserstein Distance (WD) of reconstruction and ground truth frequency distributions with the high-frequency biased support.
+For readability purposes the actual values are scaled down by a constant factor of 32768, that is the length of the sample (lower score is better).
+NERF SIREN FINER WIRE RFF KAN
+WD 184.79 218.54 205.75 131.56 362.98 79.72
+Figure 3. The comparison of spectrograms. We compare the original sample (GT) with the waveforms reconstructed using different INR approaches.:
 Spectrograms
@@
 Finally, one can note the lower average intensity in the reconstructed audio.
-Figure 3. The comparison of spectrograms. We compare the original sample (GT) with the waveforms reconstructed using different INR approaches.
-Spectral Analysis
-To demonstrate KAN’s improvements over previous INRs in high-frequency information of the reconstruction, we designed an additional experiment.
-We can treat the absolute values of the FFT of the audio signal as probability distribution of the base frequencies.
-Then, we can compare distributions of the ground truth and the reconstruction using Wasserstein distance.
-By using $[0,1,4,...(len(audio)-1)^{2}]$ as the support of the sample instead of intuitive $[0,1,2,...len(audio)-1]$, we can make the metric more biased towards high frequencies.
-Using this metric in our standard experimental setting for INRs, we obtain the following results, presented in Table 11.
-They indicate that KAN achieves the lowest score of 79.72 among all baseline methods.
-The second-best outcome is obtained by WIRE, while RFF performs the worst.
-Table 11.
-Wasserstein Distance (WD) of reconstruction and ground truth frequency distributions with the high-frequency biased support.
-For readability purposes the actual values are scaled down by a constant factor of 32768, that is the length of the sample (lower score is better).
-NERF SIREN FINER WIRE RFF KAN
-WD 184.79 218.54 205.75 131.56 362.98 79.72
 5. Conclusion
@@
 Lastly, enhancing compatibility with diverse audio encoding systems could boost reconstruction quality and expand KANs’ practical utility.
+Acknowledgements.
+This work was funded by several institutions. P. Marszałek was supported by the National Center of Science (Poland) Grant No.
+2023/50/E/ST6/00169.
+P.
+Kawa and P.
+Syga received support from the Department of Artificial Intelligence, Wrocław University of Science and Technology.
+P.
+Spurek and M.
+Rut were supported by the National Center of Science (Poland) Grant No.
+2023/50/E/ST6/00068.
 Appendix A Additional experiments
@@
 Encoding #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+Length $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 2 23,016 3.6e-3 30.46 3.31 4.26 1.21 0.66 0.42
@@
 24 52,584 3e-4 41.93 1.60 17.18 2.84 0.98 0.34
-Table 13 shows how sizes and depths of hidden layers influence results on 3 s samples.
-Similarly to 1.5 s (Table 8) the best performance is achieved by the largest network composed of [96, 48, 24] layers.
-The performance, however, does not improve with the increasing size as gradually as for 1.5 s samples.
 Table 13. KAN results on 3 s samples in relation to different number and sizes of hidden layers.
 Layer Sizes #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 [24, 12, 6] 10,500 7e-4 37.57 1.49 12.66 1.72 0.89 0.32
@@
 [96, 48, 24] 102,480 5e-4 40.96 1.18 16.12 3.52 0.99 0.16
-In comparison to the 1.5 s case, where a spline order of size 3 is clearly the optimal hyperparameter setting, its advantage is less evident in Table 14.
-KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI metrics.
+Table 13 shows how sizes and depths of hidden layers influence results on 3 s samples.
+Similarly to 1.5 s (Table 8) the best performance is achieved by the largest network composed of [96, 48, 24] layers.
+The performance, however, does not improve with the increasing size as gradually as for 1.5 s samples.
+Compared to the 1.5 s case, where a spline order of 3 is clearly optimal, its advantage is less evident in Tab.
+14.
+KAN with a spline order of 5 attains higher PSNR and SI-SNR, while remaining competitive in PESQ and STOI.
 Table 14. KAN effectiveness in terms of diverse spline order sizes and 3 s long input
 Spline #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+Order $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 1 28,860 6e-4 38.50 1.47 13.61 2.29 0.95 0.33
@@
 7 42,180 6e-4 38.88 1.32 13.97 2.20 0.96 0.34
-The results of the grid size parameter (displayed in Table 15) provide similar conclusions and tendencies as in the 1.5 s scenario (Table 10) — increasing the grid size leads to the improvement in all metrics but LSD.
+The grid size results (cf. Tab. 15) show trends similar to the 1.5 s case (Tab. 10), increasing grid size improves all metrics except LSD.
 Table 15. KAN effectiveness in terms of diverse grid size and 3 s long input
 Grid size #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 1 11,100 2.9e-3 31.34 2.00 5.46 1.38 0.77 0.40
@@
 17 46,620 6e-4 39.73 1.32 14.81 3.07 0.97 0.27
+At the end of this paragraph, we include ablation studies for the remaining hyperparameters.
+Table 16 shows the comparison of learning rates.
+A value of 0.001 yields the worst scores across all entries in the table, with the exception of the CDPAM metric.
+A learning rate of 0.01 achieves the best results in the PSNR and SI-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores.
+Generally, KAN performs well and stably with higher learning rate setting.
+It can be considered its advantage over networks like SIREN, which are sensitive to the choice of the learning rate and their learning process can collapse when too high values are used.
 Table 16. Effect of various learning rate values on 1.5 s and 3 s recordings.
 Learning #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+Rate $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 Sample Length: 1.5 s
@@
 0.001 31,080 7e-4 37.76 1.47 12.80 1.93 0.93 0.30
-At the end of this paragraph, we include ablation studies for the remaining hyperparameters.
-Table 16 shows the comparison of learning rates.
-A value of 0.001 yields the worst scores across all entries in the table, with the exception of the CDPAM metric.
-A learning rate of 0.01 achieves the best results in the PSNR and SI-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores.
-Generally, KAN performs well and stably with higher learning rate setting.
-It can be considered its advantage over networks like SIREN, which are sensitive to the choice of the learning rate and their learning process can collapse when too high values are used.
 Table 17 shows the comparison of performance of KAN depending on whether the enable scale spline flag is set to False or True.
@@
 Scale #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
-[…]
+Spline $\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
 Sample Length: 1.5 s
@@
 INR (# params) MSE PSNR LSD SI-SNR
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$
 NERF (33,261) 3.6e-3 $\pm$ 4.5e-3 32.70 $\pm$ 5.12 1.21 $\pm$ 0.42 13.16 $\pm$ 5.05
@@
 INR (# params) MSE PSNR LSD SI-SNR
-[…]
+$\rightarrow 0$ $\rightarrow\infty$ $\rightarrow 0$ $\rightarrow 100$
 NERF (33,261) 3e-3 $\pm$ 4.9e-3 34.23 $\pm$ 5.35 1.25 $\pm$ 0.52 14.81 $\pm$ 6.33